### v3

# No Nowhere Constant Continuous Function Maps All Non-Normal Numbers to Normal Numbers

Abstract
In this work, we consider the set of non-normal numbers $\mathcal{N}^{c}$ and ask if there is a non-empty open interval $I$ and a nowhere constant continuous function $\varphi:I\to\mathbb{R}$ which maps all non-normal numbers to normal numbers, i.e., $\varphi(I\cap\mathcal{N}^{c})\subset\mathcal{N}.$ We answer this question negatively. We call a set with this property super-dense. The super-density of $\mathcal{N}^{c}$ illustrates further its topological size or geometric richness. Surprisingly, the measure-theoretically “larger” set of normal numbers $\mathcal{N}$ does not share this property and we give an explicit Cantor-type construction of a nowhere constant continuous function $\hat{C}$, which maps all normal numbers to non-normal numbers. Our final result discusses the relation between super-density and the Baire categories employing various set-theoretic strategies.
Keywords— Normal numbers, Hausdorff dimension, Cantor function
1 Introduction and Main Result
Ever since Émile Borel introduced the concept of normality and showed that almost every number is normal [7], generations of mathematicians have worked hard to unravel the secrets of normal numbers. Although many profound statements about the set of normal numbers have been shown [9, 8, 13, 18, 20, 19, 15] and in recent years there has been increasing interest in the computability of concrete normal numbers [1, 4, 5], the most important conjectures concerning normal numbers remain unsolved. In particular, it currently seems unfeasible to show that important mathematical constants like $\sqrt{2},e,$ or $\pi$ are normal.
In contrast, the field of transcendental number theory has made significant breakthroughs. Although many conjectures such as Schanuel’s conjecture regarding transcendental numbers remain open, Hermite, Lindemann and Weierstrass proved in their seminal works the transcendence of $e$ and $\pi$ over a century ago [11, 12, 14, 22]. This disparity raises a natural question: why is it considerably harder to prove a number is normal than to prove it is transcendental? If one understands a proof of normality as a demonstration that a number is not non-normal, the answer seems to be that the set of non-normal numbers is of Hausdorff dimension 1, while the non-transcendental numbers, i.e. the algebraic numbers, are only a countable set. This note presents another perspective on the complexity of non-normal numbers: There is no nowhere constant continuous function that maps all non-normal numbers to normal numbers. This illustrates the intricate geometry of the set of non-normal numbers, potentially explaining the difficulty of proving normality.
To set the stage, we recall first some standard terminology. A real number $x\in{\mathbb{R}}$ can be uniquely expanded with respect to an integer base $b\geq 2$ in the form $x=X_{1}X_{2}\ldots X_{K}.x_{1}x_{2}\cdots$ with digits $X_{i},x_{i}\in\{0,1,\ldots,b-1\}$. This $b$-ary expansion of $x$ is unique if we agree to use a finite expansion, i.e., an expansion with only finitely many nonzero digits, whenever possible. We denote by $\Lambda_{b,d,M}(x)$ the relative fraction of the digit $d$ among the first $M$ fractional $x_{i}$ in base $b$ (see also (2.1)). A real number $x$ is called simply $\mathit{b}$-normal if $\lim_{M\to\infty}\Lambda_{b,d,M}(x)=\frac{1}{b}$, that is if all digits are asymptotically equidistributed in the $b$-ary expansion of $x$. If not just the digits but all finite words $w\in\bigcup_{k=1}^{\infty}\{0,1,\ldots,b-1\}^{k}$ appear with asymptotic frequency $b^{-|w|}$ in the $b$-ary expansion of $x$, we call $x$ a $\mathit{b}$-normal number. Equivalently, a real number $x$ is $b$-normal if and only if $x$ is simultaneously simply $b^{N}$-normal for all natural numbers $N\geq 1$. Finally, $x$ is an absolutely normal number or just a normal number if $x$ is $b$-normal with respect to all integer bases $b\geq 2$. Absolute normality is equivalent to the seemingly weaker condition that a number $x$ is simultaneously simply $b$-normal for all bases $b\geq 2$.
We denote by $\mathcal{N}_{b,s}=\{x\in{\mathbb{R}}\,|\,x\text{ is simply }b\text{-normal}\}$ the set of all $b$-simply normal number for some fixed integer base $b\geq 2$. Analogously, we introduce the set of all $b$-normal real numbers $\mathcal{N}_{b}$, and the set of all (absolutely) normal numbers $\mathcal{N}$. Finally, the complement of a set $A\subset{\mathbb{R}}$ is denoted by $A^{c}:={\mathbb{R}}\setminus A$.
Our results also need the concept of nowhere constant functions:
Definition 1.1.
Let $I\subset{\mathbb{R}}$ be a non-empty open interval and $\varphi:I\to{\mathbb{R}}$ some function. We say that $\varphi$ is non-constant at $t_{0}\in I$ if for every $\varepsilon>0$
$\varphi((t_{0}-\varepsilon,t_{0}+\varepsilon))\setminus\{\varphi(t_{0})\}\neq\emptyset.$ (1.1)
The function $\varphi$ is nowhere constant (on $I$) if it is non-constant at every $t\in I.$
Note that for a function $\varphi$ to be nowhere constant on some interval $I$, it suffices that the set
$D:=\left\{x\in I\,\bigg|\,\liminf_{h\to 0}\left|\frac{\varphi(x+h)-\varphi(x)}{h}\right|>0\right\}$
is dense in $I$. In particular, all typical realizations of Brownian motions are nowhere constant continuous functions.
Our first main result reads as follows:
Theorem 1.2.
Let $I\subset{\mathbb{R}}$ be any non-empty open interval and $\varphi\colon I\to{\mathbb{R}}$ a nowhere constant continuous function. Then, for every base $b\geq 2$
$\varphi(I\cap\mathcal{N}_{b,s}^{c})\cap\mathcal{N}_{b,s}^{c}\neq\emptyset.$ (1.2)
In fact, we have more generally for every countable collection $(\varphi_{k})_{k\in{\mathbb{N}}}$ of nowhere continuous functions $\varphi_{k}\colon I\to{\mathbb{R}}$,
$\bigcap_{k=1}^{\infty}\varphi_{k}(I\cap\mathcal{N}_{b,s}^{c})\neq\emptyset.$ (1.3)
Of course, (1.2) directly yields
$\varphi(I\cap\mathcal{N}^{c})\cap\mathcal{N}^{c}\neq\emptyset,$
which motivates the title of this note. A few remarks are in order.
  1. 1.
    The proof of Theorem 1.2 is spelled out in Section 2.1 for a single function $\varphi$ and its full generality in Section 2.3. Our method is constructive in the sense that under a few regularity assumptions on the $\varphi_{k}$ we can infer from the proof an algorithm computing inductively the digits of a real number $x$ such that all numbers $\varphi_{k}(x)$ are simultaneously non-normal. The details are presented in Corollary 2.9 below. Theorem 1.2 gives a (nearly) optimal result on joint non-normality, which complements the recent work [15], where the construction of simultaneously normal numbers is addressed. In the future, we hope to address simultaneous normality and non-normality, that is, to construct for example a normal number $x$ such that $x^{2}$ is not normal - mimicking the conjectured property of $\sqrt{2}$.
  2. 2.
    Let us denote by $\bar{\mathbb{Q}}$ the set of real algebraic numbers. Applying Theorem 1.2 to the countable collection $(e^{\alpha x})_{\alpha\in\bar{\mathbb{Q}}\setminus\{0\}}$ and to all open intervals $I=(\alpha,\beta)\subset{\mathbb{R}}$, one concludes that the set of numbers $x$ and all $e^{\alpha x}$ are not normal is a dense subset of $\mathbb{R}$}. Note the remarkable difference compared to the case of algebraicity: according to the celebrated Lindemann-Weierstrass Theorem $e^{x}$ is transcendental for every algebraic $x\neq 0$. While the existence of such numbers is also a consequence of the set of non-normal numbers being of second category [2, 16, 17], as far as we know there has been no constructive method to generate a non-normal number $x\neq 0$ for which $e^{x}$ is not normal. Our proof, however, can easily be turned to an algorithm for the given function $e^{x}$.
  3. 3.
    In the setting of a single nowhere constant function $\varphi$, (1.2) can be strengthened further:
    $\varphi\left(I\cap\bigcap_{b=2}^{\infty}\mathcal{N}_{b,s}^{c}\right)\cap\left(\bigcap_{b=2}^{\infty}\mathcal{N}_{b,s}^{c}\right)\neq\emptyset.$ (1.4)
    This result can be derived by an adaption of our proof strategy of Theorem 1.2 and can be found below as Corollary 2.10.
  4. 4.
    Equation (1.2) might even hold under the weaker condition that $\varphi$ is just a continuous non-constant function. However, our proof fails in this situation and it is unclear to us if the statement of Theorem 1.4 below still holds for non-constant (but not necessarily nowhere constant) functions. We hope to address this question in future research.
In our second main result, we want to argue that Theorem 1.2 should not be seen as a reflection of the measure-theoretic size of $\mathcal{N}^{c}$, but rather as a facet of the topological size and the complicated geometry of $\mathcal{N}^{c}$. To this end, it is convenient to introduce the following concept.
Definition 1.3.
We call a set $A\subset{\mathbb{R}}$ super-dense, if for every non-empty open interval $I$ and every nowhere constant continuous function $\varphi:I\to{\mathbb{R}}$ one has
$\varphi(I\cap A)\cap A\neq\emptyset.$ (1.5)
The concept of super-density appears to be novel. It is very vaguely related to the definition of a universally meager set, which is a subset $A\subset X$ of a topological space $X$ such that for every Baire space $Y$ and every nowhere constant continuous map $\varphi:Y\to X$ the preimage $\varphi^{-1}(A)$ is meager in $Y$. [21]. Clearly, a super-dense set is dense in the usual sense with respect to the Euclidean metric; and it is not hard to see that a super-dense set is necessarily uncountable. More will be said in Theorem 1.5 below. Before, we state our second result on (non-)normal numbers.
Theorem 1.4.
For every integer $b\geq 2$, there exists a super-dense set $\mathcal{Z}_{b}\subset\mathcal{N}_{b,s}^{c}$ of Hausdorff dimension zero. On the other hand, the set of absolutely normal numbers $\mathcal{N}$ is not super-dense.
The proof of Theorem 1.4 is presented in Section 2.1, where we also give an explicit definition of $\mathcal{Z}_{b}$ in (2.2). We note that Theorem 1.4 implies (1.2) (i.e. Theorem 1.2 for a single function). Due to Borel’s theorem, it might come as a surprise that the non-normal numbers are in some sense “denser” or “richer” than the absolutely normal numbers $\mathcal{N}$, which form a set of full measure. Our interpretation is that the geometry of $\mathcal{N}$ is simpler as $\mathcal{N}$ only contains numbers which behave generically, whereas $\mathcal{N}^{c}$ encodes many fractal-like subsets.
A further intuition for Theorem 1.4 can be found in the fact that $\mathcal{N}$ is a meager set or of first category [2, 16, 17]. In the following theorem we clarify the relation between super-density and Baire categories.
Theorem 1.5.
  1. 1.
    Let $A\subset\mathbb{R}$ be super-dense. Then, $A$ is also non-meager.
  2. 2.
    Let $A\subset\mathbb{R}$ be a co-meager $G_{\delta}$-set, i.e., a countable intersection of open sets. Then, $A$ is super-dense.
  3. 3.
    There exists a set $A\subset\mathbb{R}$ such that both $A,A^{c}$ are dense, but not super-dense.
  4. 4.
    There exists a set $A\subset\mathbb{R}$ such that both $A,A^{c}$ are super-dense.
The proof of Theorem 1.5 is spelled out in Section 2.2. Note that the first two assertions of Theorem 1.5 directly yield Theorem 1.4 (see also the discussion after Proposition 2.5). This more elegant approach has the disadvantage that it is not constructive and does not give an example of a nowhere constant function $\varphi$ such that $\varphi(\mathcal{N})\cap\mathcal{N}=\emptyset$. We further observe that due to Baire’s category theorem at least one of $A,A^{c}$ as in the third assertion must be non-meager, dense, but not super-dense. Of course, $\mathcal{N}$ is an example for a set which is not super-dense (and meager), and the non-normal numbers $\mathcal{N}^{c}$ are super-dense and non-meager (in fact co-meager). Finally, the sets $A,A^{c}$ from 4. are super-dense, non-meager, but not co-meager.
2 Proof of Main Results
2.1 Proof of Theorem 1.4
For the proof of Theorem 1.4, some further notation is useful. Let $b\geq 2$ and $x=0.x_{1}x_{2}x_{3}\ldots$ be the $b$-ary expansion of some $x\in[0,1)$. For $d\in\{0,1,\ldots,b-1\}$ and $M\in{\mathbb{N}}$, we set
$\Lambda_{b,d,M}(x):=\frac{1}{M}\sum_{j=1}^{M}\chi_{d}(x_{j})$ (2.1)
with the characteristic function $\chi_{d}$, that is, $\chi_{d}(d)=1$ and $\chi_{d}(k)=0$ for $k\neq d$. In other words, $\Lambda_{b,M}(x)$ is the density of the digit $d$ in the first $M$ digits of the $b$-ary expansion of $x$. For general $x\in{\mathbb{R}}$, we simply set $\Lambda_{b,d,M}(x):=\Lambda_{b,d,M}(|x|\,\mathrm{mod}\,1).$ We will be interested in the set $\mathcal{Z}_{b}$ of numbers with accumulations of zeros, that is,
$\mathcal{Z}_{b}:=\{x\in{\mathbb{R}}\,|\,\limsup_{M\to\infty}\Lambda_{b,0,M}(x)=1\}$ (2.2)
We demonstrate first that the Hausdorff dimension $\dim_{\mathcal{H}}(\mathcal{Z}_{b})=0$ as claimed in Theorem 1.4.
Lemma 2.1.
Let $b\geq 2$ be an integer. Then,
$\dim_{\mathcal{H}}(\mathcal{Z}_{b})=0.$ (2.3)
The proof of Lemma 2.1 relies on a standard combinatorial estimate and basic properties of Hausdorff measures. As it does not directly follow from Egglestone’s result on the Hausdorff-dimension of sets with specific asymptotic digit frequencies [10], we choose to sketch the complete argument in the appendix.
Proposition 2.2 below demonstrates that $\mathcal{Z}_{b}$ is super-dense and, thus, establishes the first part of Theorem 1.4.
Proposition 2.2.
For every base $b\geq 2$, the set $\mathcal{Z}_{b}$ is super-dense, that is, for every non-empty interval $I\subset{\mathbb{R}}$ and every nowhere constant continuous function $\varphi\colon I\to{\mathbb{R}}$,
$\varphi(I\cap\mathcal{Z}_{b})\cap\mathcal{Z}_{b}\neq\emptyset.$ (2.4)
The proof of Proposition 2.2 relies on an iterative “zig-zag” construction of an $x\in\mathcal{Z}_{b}$ such that $y=\varphi(x)\in\mathcal{Z}_{b}$, too. Through a two-step update we ensure that the $b$-ary expansion of $x$ and $y$ are dominated by zeros infinitely often.
Proof.
We fix the basis $b\geq 2$ and assume without loss of generality that $I\subset(0,1)$. For an $x\in(0,1)$, we denote by $x=0.d_{1}d_{2}\ldots$ its $b$-ary expansion in digits $d_{i}$. Let us introduce the set $T$,
$T:=\{x\in(0,1)\,|\,\exists L\in{\mathbb{N}}\text{ s.t. }x=0.d_{1}d_{2}\ldots d_{L}\},$ (2.5)
of numbers with finite $b$-ary expansion and we set
$\tau(x):=\inf\{L\in{\mathbb{N}}\,|\,x=0.x_{1}\ldots x_{L}\}$ (2.6)
for $x\in T$ and otherwise $\tau(x)=\infty$. Our construction a second set
$T^{\prime}:=\bigcup_{x\in T}\bigcup_{K>\tau(x)^{2}+1}\left\{x+\frac{1}{b^{K}}\right\},$ (2.7)
where the inner union runs over all integers $K>\tau(x)^{2}+1$. A number $y\in T^{\prime}$ in the form $x+b^{-K}$ with $x\in T$ and $K>\tau(x)^{2}+1$ and this decomposition is in fact unique as $K$ is determined by the position of the last non-zero digit (which must an isolated ’1’). For a $y=x+\frac{1}{b^{K}}\in T^{\prime}$, we similarly set $\tau^{\prime}(y):=K.$ The function $\tau^{\prime}(y)$ is well-defined by our previous observation. One easily sees that both, $T$ and $T^{\prime}$, are dense sets in $(0,1)$. The idea behind the definition of $T^{\prime}$ is that the zero digits of some $y=x+\frac{1}{b^{K}}\in T^{\prime}$ between the positions $\tau(x)$ and $K$ are preserved under small enough subtractions.
We now construct 4 sequences
$(x_{M}^{(1)})_{M\in{\mathbb{N}}},\,(x_{M}^{(2)})_{M\in{\mathbb{N}}},\quad(y_{M}^{(1)})_{M\in{\mathbb{N}}},\,(y_{M}^{(2)})_{M\in{\mathbb{N}}}$
and prove that $x:=\lim_{M\to\infty}x_{M}^{(2)}\in\mathcal{Z}_{b},$ $y:=\lim_{M\to\infty}y_{M}^{(2)}\in\mathcal{Z}_{b},$ and $y=\varphi(x).$
Step 1: Choice of starting points and intervals
Since $T$ is a dense set, we find some $x_{1}^{(1)}\in T\cap I$ and define $L_{x,1}:=\tau(x_{1}^{(1)}).$ Let $y_{1}^{(1)}:=\varphi(x_{1}^{(1)})$ be the corresponding function value. Moreover, we introduce the intervals
$I_{1,x}=(x_{1}^{(1)},x_{1}^{(1)}+b^{-L^{2}_{x,1}-f})\quad I_{1,y}=(-\infty,\infty),$
where we choose the integer $f\in{\mathbb{N}}$ big enough s.t. the closure of $I_{1,x}$ is still contained in $I$.
To construct the two other starting values, we recall that $\varphi$ is nowhere constant, and thus we find some $z\in I_{1,x}$ such that $\varphi(z)\neq\varphi(x_{1}^{(1)}).$ Employing the intermediate value theorem and the density of $T^{\prime}$, we in fact find some $x_{1}^{(2)}\in I_{1,x}$ such that $y_{1}^{(2)}:=\varphi(x_{1}^{(2)})\in T^{\prime}$. Note that the digits $x_{1}^{(2)}$ at position $L_{x,1}$ to $L_{x,1}^{2}$-th are zero.
Step 2: Iterative procedure:
Suppose that for some $M\in{\mathbb{N}}$, we have already defined the values and intervals
$x_{M}^{(1)}\in T,\,x_{M}^{(2)}\in I_{M,x},\quad y_{M}^{(1)}=\varphi(x_{M}^{(1)}),\,y_{M}^{(2)}=\varphi(x_{M}^{(2)})\in T^{\prime}\cap I_{M,y}.$
We describe how to construct the $M+1$-th values and intervals. We set $L_{y,M}:=\tau^{\prime}(y_{M}^{(2)})$ and choose $K_{y,M}>L_{y,M}$ big enough s.t.
$I_{M+1,y}:=\left(y_{M}^{(2)}-\frac{1}{b^{K_{y,M}}},\,y_{M}^{(2)}+\frac{1}{b^{K_{y,M}}}\right)\subset I_{M,y}.$
The $b$-ary expansion of any $z\in I_{M+1,y}$ agrees with the one of $y_{M}^{(2)}$ up to the index $L_{y,M}-1$ (here we make use of the stability of $T^{\prime}$ under subtraction due to the additional ’1’). By continuity of $\varphi$ we find some $x_{M+1}^{(1)}\in T\cap I_{M,x}$ with $y_{M+1}^{(1)}:=\varphi(x_{M+1}^{(1)})\in I_{M+1,y}.$ We set $L_{x,M+1}:=\tau(x_{M+1}^{(1)})$ and
$I_{M+1,x}:=(x_{M+1}^{(1)},x_{M+1}^{(1)}+b^{-L^{2}_{x,M+1}-1})\cap I_{M,x}\neq\emptyset$
[…]
To finish the proof, we observe that both $x_{M}^{(2)}$ and $y_{M}^{(2)}$ are Cauchy sequences and, thus, posses limits $x\in I$ and $y\in{\mathbb{R}}$. By continuity of $\varphi$, one has $y=\varphi(x).$ By construction of $x$ one immediately obtains that
$\Lambda_{b,0,L_{M,x}^{2}}(x)\geq 1-\frac{1}{L_{M,x}},$
which implies that $x\in\mathcal{Z}_{b}$ if $L_{M,x}\to\infty.$ If $L_{M,x}$ remains bounded the limit $x$ has only finitely many non-zero digits and thus $x\in\mathcal{Z}_{b},$ too. One similarly shows that $y\in\mathcal{Z}_{b},$ which completes the proof. ∎
It remains to show that the set of absolutely normal numbers $\mathcal{N}$ is not super-dense. In fact, we will construct an even larger set $\mathcal{A}\supset\mathcal{N}$, which is not super-dense. We define
[…]
and recall that the set of $2$-simply normal numbers is given by
$\mathcal{N}_{s,2}=\left\{x\in{\mathbb{R}}\,|\,\lim_{M\to\infty}\Lambda_{2,0,M}(x)=\frac{1}{2}\right\}.$
We set
$\mathcal{A}:=\mathcal{T}_{1}\cap\mathcal{N}_{s,2}.$ (2.8)
Of course, $\mathcal{N}\subset\mathcal{A}$. Hence, $\mathcal{A}^{c}$ is a null set. We want to show
Proposition 2.3.
The set $\mathcal{A}$ is not super-dense.
Proof.
The proof heavily relies on properties of Cantor’s function $C:[0,1)\to[0,1),$ also commonly known as devil’s staircase. We recall that $C$ is increasing and continuous. For our purposes, the following description of $C$ is handy. Let $x=0.x_{1}x_{2}x_{3}\ldots$ be the ternary expansion of $x\in[0,1)$. If the ternary expansion of $x$ contains no ’1’, one sets $C(x):=x$. Otherwise, there exists a minimal index $m\in{\mathbb{N}}$ s.t. $x_{m}=1$ and we set $y=C(x)$ to be the number whose binary(!) expansion $y=0.y_{1}y_{2}y_{3}\cdots$ is given by
$y_{i}:=\begin{cases}0,&\text{ if }i>m,\\ \min\{x_{i},1\},&\text{ if }i\leq m.\end{cases}$
Note that $C$ maps all numbers in $\mathcal{T}_{1}$ to values with finite binary expansion. That is,
$C(\mathcal{A}\cap(0,1))\cap\mathcal{A}=\emptyset.$ (2.9)
Unfortunately, $C$ is locally constant at the deleted thirds in the iterative construction of the Cantor set. The main idea is to enrich $C$ by self-similar copies of itself in each interval such that it becomes a nowhere constant function, which is continuous and still satisfies (2.9).
To this end, it is convenient to work with the symmetrized version $C_{s}$
$C_{s}(x):=\begin{cases}C(x),&\text{ if }x\leq\frac{1}{2},\\ C(1-x),&\text{ if }x>\frac{1}{2},\end{cases}$ (2.10)
which still is continuous and preserves the crucial property of the Cantor function that it maps all numbers with at least a ’1’ in the ternary expansion to a number with finite binary expansion. We now define the enriched Cantor function $\hat{C}:[0,1]\to[0,1]$. To this end, we distinguish the cases, where $x$ has finitely many ’1’s and infinitely many. If $x$ has no ’1’, we set $\hat{C}(x)=C_{s}(x).$ If $x$ has $m$ ’1’s in its ternary expansion, we write $x=0.w_{1}w_{2}\ldots w_{m}\sigma,$ where the finite words $w_{i}$ contain exactly an ’1’ as their last digit - we call such words one-stripped in this proof - and $\sigma$ is a word without any ’1’. Then, we set
$\hat{C}(x):=\left(\sum_{r=1}^{m}2^{-r\sum_{k=1}^{r-1}|w_{r}|}C_{s}(w_{r})\right)+2^{-(m+1)\sum_{k=1}^{r}|w_{r}|}C_{s}(\sigma),$ (2.11)
where the empty sum is understood as zero. Here, we abuse notation a bit as we identify a ternary word $w$ with the real number $x=0.w$, i.e. $C_{s}(w)$ should be understood as $C_{s}(x)$ with $x=0.w$. In the case $m=1$, one applies the normal Cantor function to the first part $w_{1}$, which gives a number terminating after $|w_{1}|$ bits. Then, one inserts $|w_{1}|$-many zeros and continues with $\sigma$ - reinterpreted as bit strings. For higher $m$, one alternates between inserting an increasing number of zeros and inserting the Cantor bit-string of each $w_{i}$ afterwards. If $x$ has infinitely many ’1’s in its ternary expansion, we write $x=0.w_{1}w_{2}\ldots$ with the same definition for $w_{i}$ as above and one similarly sets
$\hat{C}(x):=\sum_{r=1}^{\infty}2^{-r\sum_{k=1}^{r-1}|w_{r}|}C_{s}(w_{r}).$ (2.12)
The function $\hat{C}(x)$ is nowhere constant as $C(x)\neq C(y)$ for any $y$ with finite ternary expansion and any $x$ with infinitely many ’0’s and ’2’s.
Next, we verify that $\hat{C}(x)$ is continuous. We consider the approximations $\hat{C}_{k}(x),$ which roughly speaking stop after the $k$-th ’1’ (similarly to the original Cantor function which stops after the first ’1’). A rigorous construction goes as follows. Let $T_{k}:[0,1)\to[0,1)$ be the truncation map
[…]
where the $w_{i}$ are the unique one-stripped words from above with a single ’1’ as their last digit. We simply set $\hat{C}_{k}(x):=\hat{C}(T_{k}(x)).$ Let us show that $\hat{C}_{k}(x)$ is continuous. We note first that $\hat{C}_{1}(x)=C_{s}(x)$ and for each natural number $k$ the difference $\hat{C}_{k+1}-\hat{C}_{k}$ is nonzero only on closed intervals of the form
$I_{w_{1},w_{2},\ldots,w_{k}}:=[0.w_{1},w_{2},\ldots,w_{k},w_{1},w_{2},\ldots,w_{k}+3^{-\sum_{j=1}^{k}w_{j}}]$
with one-stripped $w_{i}$ and on each such interval we have
$\hat{C}_{k+1}(0.w_{1}\ldots w_{k}\tau)-\hat{C}_{k+1}(0.w_{1}\ldots w_{k}\tau)=2^{-(k+1)\sum_{j=1}^{r}|w_{j}|}C_{s}(\tau)$
for any ternary word $\tau$. The continuity is evident as one inserts a damped copy of the original (symmetrized) Cantor function in each of the separated intervals $I_{w_{1},w_{2},\ldots,w_{k}}$ , where $\hat{C}_{k}$ is constant. Moreover, by construction
$\sup_{x\in(0,1)}|\hat{C}(x)-\hat{C}_{k}(x)|\leq 2^{-k^{2}}$
so that $\hat{C}_{k}$ converges uniformly to $\hat{C}$. This implies that $\hat{C}$ is continuous, too.
It remains to show that $\hat{C}(\mathcal{T}_{1})\cap\mathcal{N}_{2,s}=\emptyset.$ To this end, let $x\in\mathcal{T}_{1}$ and we write for its ternary expansion $x=0.w_{1}w_{2}\cdots$ with one-stripped $w_{i}$. Let $\hat{w_{i}}$ be the “Cantor image” of $w_{i}$ that is the binary string obtained by flipping each ’2’ in $w_{i}$ to a ’1’. Then, the binary expansion of $\hat{C}(x)$ has the form $\hat{C}(x)=\hat{w_{1}}z_{1}\hat{w_{2}}z_{2}\cdots$, where each $z_{r}$ is a string of zeros. Now, note that after the last digit of $z_{m}$ the binary expansion of insertion of $\hat{C}(x)$ has at most $\sum_{r=1}^{m}w_{r}$ ’1’s but at least $(m-1)\sum_{r=1}^{m}w_{r}$ zeros. That is there a sequence of integers $L_{m}$ s.t.
$\Lambda_{2,0,L_{m}}(C(x))\geq 1-\frac{1}{m}.$
This finishes the proof. ∎
2.2 Proof of Theorem 1.5
In this section, we prove all statements of Theorem 1.5 one by one. The proofs are all independent of each other and are therefore presented separately. We start with the first statement.
Proposition 2.4.
Let $A\subset\mathbb{R}$ be super-dense. Then, $A$ is also non-meager.
Since the normal numbers $\mathcal{N}$ form a meager set this result gives a further justification for the fact that $\mathcal{N}$ is not super-dense.
Proof.
We proceed by contraposition. We fix some meager set $A\subset\mathbb{R}$ and we want to show that $A$ is not superdense. Our argument is nonconstructive. Indeed, let $I=[0,1]$ be the unit interval, $C(I)$ the space of continuous function on $I$ - equipped with the uniform topology - and consider the sets
$Y$ $:=\{f\in C(I)\,|\,f\text{ is constant somewhere}\},$ (2.13)
$Z$ $:=\{f\in C(I)\,|\,f(I\cap A)\cap A\neq\emptyset\}.$ (2.14)
Since $C(I)$ is a complete metric space, it is enough to show that $Y$ and $Z$ are meager as then Baire’s category theorem implies that there exist a nowhere constant $f\in C(I)$ such that $f(I\cap A)\cap A=\emptyset$. Extending such a $f$ linearly on the whole real line, proves that $A$ is not super-dense.
Step 1: $Y$ is meager.
If a function $f$ is constant at some $t$, there exists an interval $q_{1},q_{2}]$ with rational endpoints on which $f$ is constant. Thus,
[…]
It suffices to show that each $Y_{q_{1},q_{2}}$ is nowhere dense. Clearly, $Y_{q_{1},q_{2}}$ is a closed set in $C(I)$ and for any $f\in Y_{q_{1},q_{2}}$ and any $\varepsilon>0$, the function $g_{\epsilon}(x)=f(x)+\varepsilon x\notin Y_{q_{1},q_{2}}$. Hence, $Y_{q_{1},q_{2}}$ is nowhere dense.
Step 2: $Z$ is nowhere dense if $A$ is nowhere dense
Suppose that $A\subset\mathbb{R}$ is closed and nowhere dense. We show that then $Z$ defined as in (2.14) is nowhere dense (in $C(I)$), too. We fix some $f\in C(I)$ and $\varepsilon>0$. Since $A$ is nowhere dense, we find for every $x\in I$ a $u_{x}\in A^{c}$ such that $|u_{x}-f(x)|<\varepsilon$. Due to the continuity of $f$, the sets
$U_{x}:=\{y\in I\,|\,|f(y)-u_{x}|<\varepsilon\}$
are (relatively) open and let $\hat{U}_{x}$ further be the connected component of $U_{x}$ containing $x$. Then, the collection $(\hat{U}_{x})_{x\in I}$ forms an open cover of $I$. So, by compactness, there exists a finite cover $\hat{U}_{x_{1}},\ldots\hat{U}_{x_{n}}$ of $I$ with $x_{1}<x_{2}<\cdots<x_{n}$. Using again the assumption that $A$ is closed and nowhere dense, we can find open intervals $J_{x_{i},x_{i+1}}\subset(x_{i},x_{i+1}$ such that $J_{x_{i},x_{i+1}}\subset A^{c}$ for $i=1,\ldots,n-1$. Finally, we set $I_{x_{i}}:=\hat{U}_{x_{i}}\setminus\left(\bigcup_{j=1}^{n-1}J_{x_{j},x_{j+1}}\right)$ for $i=1,\ldots,n$. The intervals $I_{x_{i}},J_{x_{i},x_{i+1}}$ form a partition of $I$ and we define the function $g$ to be $g\equiv u_{x_{i}}$ on $I_{x_{i}}$ and the linear interpolation between $u_{x_{i}}$ and $u_{x_{i+1}}$ on $J_{x_{i},x_{i+1}}$. The function $g$ is clearly continuous and by construction $\|f-g\|<\varepsilon$ and $g(A\cap I)\cap A=\emptyset$. Since $\varepsilon>0$ is arbitrary, $Z$ is nowhere dense.
Step 3: $Z$ is meager.
Since $A$ is meager, we find a countable family of nowhere dense sets $F_{k}\subset\mathbb{R}$ such that $A=\bigcup_{k=1}^{\infty}F_{k}$. For $K\in\mathbb{N}$, we define $G_{K}:=\bigcup_{k=1}^{K}\bar{F}_{k}$ and
$Z_{K}:=\{f\in C(I)\,|\,f(I\cap G_{K})\cap G_{K}\neq\emptyset\}.$
We clearly have $Z\subset\bigcup_{K=1}^{\infty}Z_{K}$ and by the second step each $Z_{K}$ is nowhere dense. Hence, $Z$ is a meager subset of $C(I)$. ∎
While Proposition 2.4 gives a rather satisfactory necessary condition for a real set to be super-dense, the next result deals with a complementary sufficient criterion
Proposition 2.5.
Let $A\subset\mathbb{R}$ be a co-meager $G_{\delta}$-set. Then, $A$ is super-dense.
Let us first demonstrate that the not simply $b$-normal numbers $\mathcal{N}_{b,s}^{c}$ contain a co-meager $G_{\delta}$-set $G_{b}$. Then, by Proposition 2.5, $G_{b}$ is super-dense and, thus, $\mathcal{N}_{b,s}^{c}$ is, too. As before, for $x\in{\mathbb{R}}$ we denote by $x_{i}\in\{0,\ldots,d-1\}$ the decimal places of $x$. We define
$G_{n}:=\bigcup_{k\geq n}\{x\in{\mathbb{R}}\setminus\mathbb{Q}\,|\,x_{m}=0\text{ for all }m=k,\ldots,k^{2}\}=:\bigcup_{k\geq n}G_{n,k}$
The set $G_{n}$ is clearly dense. It is also open since each $G_{n,k}$ is a finite union of intervals with rational boundary points. Hence, $G_{n}^{c}$ is nowhere dense and
$G_{b}:=\bigcap_{n\geq 1}G_{n}$ (2.15)
is a co-meager $G_{\delta}$ set. Finally we note that $G_{b}\subset\mathcal{Z}_{b}\subset\mathcal{N}_{b,s}^{c}.$
The following proof follows the suggestion of the anonymous referee.
Proof.
Let $\varphi:\mathbb{R}\to\mathbb{R}$ be continuous and nowhere constant. By the intermediate value property, for every open interval $I$ there exists an open subinterval $I_{0}\subseteq I$ such that $\varphi(I_{0})$ contains an open set $J$ in its closure.
We now show that $\varphi(A\cap I_{0})$ is co-meager in some open interval. By hypothesis, $A$ is co-meager in $\mathbb{R}$} and, hence, $A\cap I_{0}$ is co-meager in $I_{0}$. Of course, $A\cap I_{0}$ is also dense in $I_{0}$. Since $\varphi$ is continuous with $\varphi(I_{0})\supseteq J$, the set $\varphi(A\cap I_{0})$ is dense in $J$. Since $A$ is a $G_{\delta}$ set, $\varphi(A\cap I_{0})$ is the continuous image of a Borel set and, consequently, $\varphi(A\cap I_{0})$ has the Baire property.
Then there exists an open set $U$ such that the symmetric difference $\varphi(A\cap I_{0})\triangle U$ is meager. Since $\varphi(A\cap I_{0})$ is dense in $J$, we have that $U$ is dense in $J$, so there is some open interval $J^{\prime}\subseteq J$ such that $J^{\prime}\subseteq U$ and $U\cap J^{\prime}$ is dense in $J^{\prime}$. This means $\varphi(A\cap I_{0})\cap J^{\prime}$ differs from the dense open set $U\cap J^{\prime}$ by a meager set, hence $\varphi(A\cap I_{0})\cap J^{\prime}$ is co-meager in $J^{\prime}$.
We have two sets that are co-meager in the open interval $J^{\prime}$: $\varphi(A\cap I_{0})\cap J^{\prime}$ (by the argument above) and $A\cap J^{\prime}$ (by the initial assumption that $A$ is co-meager and, thus, co-meager in every open interval). By the Baire Category Theorem, the intersection of two co-meager subsets of the complete metric space $J^{\prime}$ is non-empty. Thus,
$\varphi(A\cap I_{0})\cap A\supseteq(\varphi(A\cap I_{0})\cap J^{\prime})\cap(A\cap J^{\prime})\neq\emptyset.$
Our proofs of the remaining assertions of Theorem 1.5 are in some sense more explicit. We recall the popular Vitali construction of a non-measurable real set $V$: define the equivalence relation on $\mathbb{R}$
$x\simeq y\Leftarrow x-y\in\mathbb{Q},$
which naturally decomposes $\mathbb{R}$} into equivalence classes $[x]$. We select exactly one representative of each equivalence class to form the set $V$. Of course, we made use of the axiom of choice in this construction. It is easy to see that we can assume $V$ to be dense. Indeed let $D\subset V$ be some infinitely countable set and $(I_{q_{1},q_{2}})_{q_{1}<q_{2}\in\mathbb{Q}}$ be the countable collection of open intervals with rational endpoints $q_{1},q_{2}$. After enumerating both countable sets, we replace each $d_{k}\in D$ by some $d_{k}+r_{k}\in I_{k}$ with $r_{k}\in\mathbb{Q}$. We fix such dense “Vitali” set $V$ and define $V_{q}:=\{x+q\,|\,x\in V\}$ for $q\in\mathbb{Q}$. By construction all $V_{q}$ are pairwise disjoint. We further introduce the following partition of the real line
$E:=\bigcup_{n\in{\mathbb{Z}}}[2n,2n+1)\qquad F:=\bigcup_{n\in{\mathbb{Z}}}[2n-1,2n).$
After these preparations, we are ready to state our next result
Proposition 2.6.
Let $A:=\bigcup_{q\in E\cap\mathbb{Q}}V_{q}$. Then, $A^{c}=\bigcup_{q\in F\cap\mathbb{Q}}V_{q}$ and both $A,A^{c}$ are dense, but not super-dense.
Proof.
The representation for $A^{c}$ follows from the fact that the sets $V_{q}$ form a partition of the real line. Density is a direct consequence of $V$ itself being dense. Considering $\varphi(x):=x+1$, we directly see that $\varphi(A)=A^{c}$ and $\varphi(A^{c})=A$. Hence, neither of the sets is super-dense. ∎
For the final assertion, we need a few set theoretic preliminaries. We denote as usual by $\aleph_{0}$ the cardinality of $\mathbb{N}$} and by the $\aleph_{1}$ the smallest uncountable cardinality, i.e., the cardinality of the set of all countable ordinal numbers. The cardinality of the the reals is given by $2^{\aleph_{0}}$ and under the continuum hypothesis $2^{\aleph_{0}}=\aleph_{1}.$ We further recall that any set $X$ can be well-ordered and for such a well-order $\prec$ we set $I_{\prec}(x)=\{t\in X\,|\,t\prec x\}$.
Lemma 2.7.
  1. 1.
    The cardinality of the set of nowhere constant continuous function is
    $\{f:{\mathbb{R}}\to{\mathbb{R}}\,|\,f\text{ is continuous and nowhere constant}\}=2^{\aleph_{0}}.$
  2. 2.
    Let $X$ be a set. Then, there exists a well-order such that $|I_{\prec}(x)|<|X|$.
Proof.
  1. 1.
    Since the set contains all functions of the form $x+\alpha$ with $\alpha\in{\mathbb{R}}$, its cardinality is at least $2^{\aleph_{0}}$. On the other hand, since every continuous function is uniquely determined by its values on $\mathbb{Q}$ the cardinality is at most $|{\mathbb{R}}^{\mathbb{Q}}|=2^{\aleph_{0}}$.
  2. 2.
    This statement is folklore in set theory (see e.g. [3][Theorem 2]). Suppose a well-order $\prec$ has not the property. Then, there is a (with respect to $\prec$) minimal $y\in X$ such that $|I_{\prec}(y)|=|X|$. By definition of cardinality numbers, there exists a bijection $\tau:I_{\prec}(y)\to X$. The new well-order $x\prec^{\prime}x^{\prime}\Leftrightarrow\tau^{-1}(x)\prec\tau^{-1}(x^{\prime})$ has the desired property.
We are ready to complete the proof of Theorem 1.5.
Proposition 2.8.
There exists a set $A\subset\mathbb{R}$ such that both $A,A^{c}$ are super-dense.
Proof.
Let $F:=\{f:{\mathbb{R}}\to{\mathbb{R}}\,|\,f\text{ is continuous and nowhere constant}\}$ and $\prec$ a well-order on $F$ with the property from Lemma 2.7. Let $q_{n}$ be an enumeration of the rationals. We construct via transfinite induction disjoint sets $A_{f}^{1},A_{f}^{2}$ with $f\in F$ such that
  1. 1.
    $|A_{f}^{i}|<2^{\aleph_{0}}$ for every $f\in F$,
  2. 2.
    $A_{f}^{i}\subset A_{f^{\prime}}^{i}$ for $f\prec f^{\prime}$,
  3. 3.
    for every $g\prec f$ and every $q_{n}$ there exists an $r\in A_{f}^{i}$ such that $|r-q_{n}|<1/n$ and $g(r)\in A_{f}^{i}$.
Fix $f\in F$ and suppose we have already constructed $A_{g}^{i}$ for $g\prec f,g\neq f$. We first set
$B_{f}^{i}:=\bigcup_{g\prec f,g\neq f}A_{g}^{i},$
which by assumption satisfies the first two conditions and the third condition except for the function $f$. The $f$-update is done inductively, i.e., a standard induction on $n$, where at each step only finitely many points are added. Let $A_{f,0}^{i}:=B_{f}^{i}$ and suppose we already know $A_{f,n-1}^{i}$. Since $f$ is nowhere constant, the image $f((q_{n}-1/n,q_{n}+1/n))$ contains an interval and thus has cardinality $2^{\aleph_{0}}$. This is still true for $X=f((q_{n}-1/n,q_{n}+1/n))\setminus(\cup_{i=1,2}A^{i}_{f,n-1}\cup f(A^{i}_{f,n-1}))$. We pick $y\in X$ and $x\in(q_{n}-1/n,q_{n}+1/n)$, and set $A^{1}_{f,n}:=A^{1}_{f,n-1}\cup\{y\}\cup\{x\}$. We repeat the construction for $A^{2}_{f,n}$ and note that $A^{2}_{f,n}$$A^{1}_{f,n}$ are still disjoint and have cardinality less than $2^{\aleph_{0}}$.
Having constructed the sets $A^{i}_{f,n}$, we define $A^{i}_{f}=\bigcup_{n\in{\mathbb{N}}}A^{i}_{f,n}$ and by construction these sets have the desired properties.
We define $A:=\bigcup_{f\in F}A^{1}_{[}f]$ and due to the third property of $A_{f}^{1}$, the set $A$ is super-dense. Since $A^{c}$ contains all sets $A_{f}^{2}$, it is super-dense by the same reasoning. ∎
2.3 Proof of Theorem 1.2
The proof of Theorem 1.2 in the case of a countable collection of nowhere constant continuous functions $\varphi_{k}:I\to{\mathbb{R}}$ largely builds on the construction for a single function $\varphi$ presented in the proof of Proposition 2.2.
Proof of Theorem 1.2 .
Let $b\geq 2$ be some integer base, with loss of generality we take $I=(0,1)$ and let $(\varphi_{k})_{k\in{\mathbb{N}}}$ be a family of nowhere constant continuous functions on $I$. The construction of a number $x$ such that all $\varphi_{k}(x)$ are simultaneously non-normal is essentially the same procedure as in Proposition 2.2, but controlling all $\varphi_{k}$ makes its description more tedious and notation-heavy. We fix some map $\sigma:{\mathbb{N}}\to{\mathbb{N}}$ with the property that $|\sigma^{-1}\{k\}|=\infty$ for any $k\in{\mathbb{N}}$. A simple and computable example for such a map is given by
$\sigma(N^{2}+m):=m+1\quad\text{ for }N\in{\mathbb{N}}\text{ and }0\leq m\leq 2N.$
At step $m$ of our construction, the function $\varphi_{\sigma(m)}$ is of main interest. We denote by
$S_{m}:=\sigma^{-1}(\{1,\ldots,m\})$
the so far traversed indices and $L_{m}(k):=|\{1,\ldots,m\}\cap\sigma^{-1}(k)|$ the number of times the integer $k$ has been visited. The numbers $y^{(k)}_{j}\in T^{\prime}$ stand for the $j$-th approximation of $\varphi_{k}(x)$. We denote by $x^{(m)}$ the approximations of the desired number $x\in I$. The idea is to ensure that $y_{k}:=\varphi_{k}(x)=\lim_{j\to\infty}y^{(j)}_{k}$ and to force $y_{k}$ to be zero-heavy, i.e., $y_{k}\in\mathcal{Z}_{b}$.
Step 1: Choice of starting points
Using continuity and the nowhere constancy of $\varphi_{\sigma(1)}$, we find as in the proof of Proposition 2.2 a real $x^{(1)}\in(1/4,3/4)$ and $y_{\sigma(1)}^{(1)}\in T^{\prime}$ such that $y_{\sigma(1)}^{(1)}=\varphi_{\sigma(1)}(x^{(1)})$.
Step 2: Iterative procedure
Suppose we have already constructed $x^{(m)}\in I$ and approximants $y_{k}^{(L_{m}(k))}\in T^{\prime}$ such that
$|y_{k}^{(L_{m}(k))}-\varphi_{k}(x_{m})|<b^{-\tau^{\prime}\left(y_{k}^{(L_{m}(k)))}\right)-1}$ (2.16)
for all $k\in S_{m}$. Here we used the function $\tau^{\prime}$ from the proof of Proposition 2.2. Our goal is to find $x^{(m+1)}\in I$ and $y^{(L_{m+1}(\sigma(m+1))}_{\sigma(m+1)}\in T^{\prime}$ such that $|x^{(m+1)}-x^{(m)}|<2^{-(m+2)}$, $\tau^{\prime}\left(y^{(L_{m+1}(\sigma(m+1))}_{\sigma(m+1)}\right)>m$ and (2.16) holds true for $m+1$, too.
Denote by $U_{m}$ be the open set, for which $z\in U_{m}$ satisfies the condition (2.16) and $|z-x^{(m)}|<2^{-(m+2)}$. The standard argument involving the nowhere constancy and continuity of $\varphi_{\sigma(m+1)}$ and the density of $T^{\prime}$, we find an $x^{(m+1)}\in U_{m}$ such that $y^{(L_{m+1}(\sigma(m+1))}_{\sigma(m+1)}:=\varphi_{\sigma(m+1)}(x^{(m+1)})\in T^{\prime}$ with $\tau^{\prime}\left(y^{(L_{m+1}(\sigma(m+1))}_{\sigma(m+1)}\right)>m.$
Step 3: Conclusion
The numbers $x^{(M)}$ form a Cauchy sequence with limit $x$. Since $x^{(1)}\in(1/4,3/4)$ and $|x^{(m+1)}-x^{(m)}|<2^{-(m+2)}$, we have $x\in(0,1)$. Since for any $k\in{\mathbb{N}}$ the sequence $y_{k}^{(m)}$ forms a subsequence of the function values $\varphi_{k}(x_{m})$, we have by continuity $y_{k}:=\varphi_{k}(x)\lim_{m\to\infty}y_{k}^{(m)}$. On the other hand,
$|y_{k}^{(m+1)}-y_{k}^{(m)}|<b^{-\tau^{\prime}\left(y_{k}^{(m)}\right)-1)},$
which guarantees that there are infinitely many integers $K\in{\mathbb{N}}$ such that all digits of $y_{k}$ between positions $K$ and $K^{2}$ are zero. Hence, $y_{k}\in\mathcal{Z}_{b}$ for all $k\in{\mathbb{N}}$. ∎
We continue our discussion of Theorem 1.2 by giving a rigorous account on when the above construction of $x$ can be turned to an algorithm.
Corollary 2.9.
Suppose the functions $\varphi_{k}:(0,1)\to{\mathbb{R}}$ satisfy for every $k\in{\mathbb{N}}$ the following conditions.
  1. 1.
    Each $\varphi_{k}$ is a computable function.
  2. 2.
    $\varphi_{k}$ is a $C^{1}$-function, i.e., continuously differentiable.
  3. 3.
    The set $\{\varphi_{k}^{\prime}(x)=0\}$ is finite and its size and elements are computable.
Then, there exists an algorithm computing an $x\in(0,1)$ such that all $\varphi_{k}(x)$ are simultaneously non-normal.
The first assumption is essentially minimal and the latter two can certainly be further weakened, but are in all practical satisfied and simplify the discussion of the algorithm.
Proof.
We apply the same construction as in the proof of Theorem 1.2. Hence, it suffices to show that there are computable choices for $y_{k}^{(m)}\in T^{\prime}$ and $x^{(m)}$. We only demonstrate that the iterative update is computable. Note that due to the second assumption each $\varphi_{k}$ is piecewise strictly monotone. Moreover, since all functions $\varphi_{k}$ are computable, a bisection procedure allows us to compute an $\delta_{m}>0$ such that for all $z\in[x^{(m)}-\delta_{m},x^{(m)}+\delta_{m}]$,
$|y_{k}^{(L_{m}(k))}-\varphi_{k}(z)|<b^{-\tau^{\prime}\left(y_{k}^{(L_{m}(k)))}\right)-1}.$
Computing in a next step, the solutions to $\varphi^{\prime}_{\sigma(m+1)}(z)=0$ to a sufficient precision, we can select $z_{1},z_{2}\in[x^{(m)}-\delta_{m},x^{(m)}+\delta_{m}]$ such that $\varphi_{\sigma(m+1)}(z_{1})\neq\varphi_{\sigma(m+1)}(z_{2})$. In a next step, we compute $\varphi_{\sigma(m+1)}(z_{i})$ to precision $\varepsilon>0$ (denoted by $\varphi_{\sigma(m+1),\varepsilon}(z_{i})$) such that
$|\varphi_{\sigma(m+1),\varepsilon}(z_{1})-\varphi_{\sigma(m+1),\varepsilon}(z_{2})|>4\varepsilon.$
Choosing $y^{(L_{m+1}(\sigma(m+1))}_{\sigma(m+1)}\in T^{\prime}$ such that
$\left|y^{(L_{m+1}(\sigma(m+1))}_{\sigma(m+1)}-(\varphi_{\sigma(m+1),\varepsilon}(z_{1})+\varphi_{\sigma(m+1),\varepsilon}(z_{2}))/2\right|<\varepsilon$
we can be sure that there exists some $w\in[x^{(m)}-\delta_{m},x^{(m)}+\delta_{m}]$ such that $y^{(L_{m+1}(\sigma(m+1))}_{\sigma(m+1)}=\varphi_{\sigma(m+1),\varepsilon}(w)$. In contrast to the proof of Theorem 1.2, we cannot set $x^{(m+1)}=w$, but it is enough to find an $z$ such that (2.16) holds for $m+1$. We can find such an $x^{(m+1)}$ y a bisection procedure. ∎
Our last result deals with the non-normality for various bases.
Corollary 2.10.
Let $I\subset{\mathbb{R}}$ be any non-empty open interval and $\varphi\colon I\to{\mathbb{R}}$ a nowhere constant continuous function. Then,
$\varphi\left(I\cap\bigcap_{b=2}^{\infty}\mathcal{N}_{b,s}^{c}\right)\cap\left(\bigcap_{b=2}^{\infty}\mathcal{N}_{b,s}^{c}\right)\neq\emptyset$ (2.17)
Proof.
A constructive proof follows essentially the lines of the proof of Theorem 1.2. Instead, we give a topological proof using Theorem 1.5. Recall the co-meager $G_{\delta}$-set $G_{b}$ from (2.15). We simply define $G:=\bigcup_{b\geq 2}G_{b}$, which is still a co-meager $G_{\delta}$-set. By Theorem 1.5, $G$ is super-dense. Since $G_{b}\subset\mathcal{Z}_{b}$, we also have $G\subset\bigcap_{b=2}^{\infty}\mathcal{N}_{b,s}^{c}$, which proves the super-density of $\bigcap_{b=2}^{\infty}\mathcal{N}_{b,s}^{c}$. ∎
Appendix A Proof of Lemma 2.1
Proof of Lemma 2.1.
First, we recall the construction of the (outer) Hausdorff measure on the real line. For real numbers $s,\delta>0$ we introduce the functional
$\mathcal{H}^{s}_{\delta}(A):=\inf\left\{\sum_{k}^{\infty}|I_{k}|^{s}\,\bigg|\,A\subset\bigcup_{k=1}^{\infty}I_{k}\text{ and }(I_{k})_{k}\text{ are open intervals with }|I_{k}|<\delta\right\}$ (A.1)
for any set $A\subset{\mathbb{R}}$. Then, the $s$-dimensional (outer) Hausdorff is defined as
$\mathcal{H}^{s}(A):=\sup_{\delta>0}\mathcal{H}^{s}_{\delta}(A)=\lim_{\delta\to 0}\mathcal{H}^{s}_{\delta}(A).$
A set $A\subset{\mathbb{R}}$ is of zero Hausdorff dimension if and only if $\mathcal{H}^{s}(A)=0$ for all $s>0.$ Due to subadditivity of the Hausdorff measure, it suffices to show
$\mathcal{H}^{s}_{\delta}(\mathcal{Z}_{b}\cap[0,1))$ (A.2)
for all $s,\delta>0$. From now on, we fix the integer $b\geq 2$ and $s,\delta$. We proceed in two steps.
Step 1: Combinatorial bound on zero-heavy finite words
Let us fix some $0<\varepsilon<\frac{1}{2}$ and for $M\in{\mathbb{N}}$ we consider the set $\Omega_{M}:=\{0,1,\ldots,b-1\}^{M}$ of $b$-ary words of length $M$. For any $w\in\Omega_{M}$ we set
$N_{M}(w)=\frac{1}{M}\sum_{j=1}^{M}\chi_{0}(w_{j}),$
and define the set
$\Omega_{M}^{\varepsilon}:=\{w\in\Omega_{M}\,|\,N_{m}(w)\geq 1-\varepsilon\}.$
We note that the size of $\Omega_{M}^{\varepsilon}$ can be written in terms of a binomial expansion
$|\Omega_{M}^{\varepsilon}|=\sum_{K\geq(1-\varepsilon)M}^{M}\binom{M}{K}(b-1)^{M-K}\leq\varepsilon M\binom{M}{\lceil(1-\varepsilon)M\rceil}(1-b)^{\varepsilon N},$ (A.3)
where the bound follows from the monotonicity of the binomial. We now employ Stirling’s formula to obtain the asymptotic expansion
$\binom{M}{\lceil(1-\varepsilon)M\rceil}=\exp((\gamma(\varepsilon)+o(1))M),$
where we used Landau’s $o$-notation and introduced the binary entropy $\gamma:[0,1]\to{\mathbb{R}},$
$\gamma(x)=-x\ln(x)-(1-x)\ln(1-x).$ (A.4)
Note that by the above we may find some $M_{0}=M_{0}(\varepsilon)$ s.t. that for all $M\geq M_{0}(\varepsilon)$
$|\Omega_{M}^{\varepsilon}|\leq\exp((2\gamma(\varepsilon)+\varepsilon\ln(1-b))M)$ (A.5)
and that the prefactor in the exponential tends to zero as $\varepsilon\to 0$. Therefore, for any $s>0$, there exists some $\varepsilon(s)$ s.t. for all $\varepsilon<\varepsilon(s)$
$2\gamma(\varepsilon(s))+\varepsilon(s)\ln(1-b)<\frac{s\ln(b)}{2}$ (A.6)
Step 2: Construction of a suitable cover
We denote by $\phi_{M}:[0,1)\to\Omega_{M}$ the function, which maps $x$ to the word $w$ consisting of the $M$ first digits of $x$ with respect to $b$. The partial inverse $\psi_{M}:\Omega_{M}\to[0,1)$ maps a word $w$ to the number $x=0.w_{1}w_{2}\ldots w_{M}.$ By definition,
$\mathcal{Z}_{b}\cap[0,1)\subset\bigcup_{M=K}^{\infty}\{x\,|\,N_{M}(x)\geq 1-\varepsilon\}=\bigcup_{M=K}^{\infty}\{x\,|\,\phi_{M}(x)\in\Omega_{M}^{\varepsilon}\}$ (A.7)
for any integer $K$ and any $\varepsilon>0.$ Moreover, we note that
$\{x\,|\,\phi_{M}(x)\in\Omega_{M}^{\varepsilon}\}\subset\bigcup_{w\in\Omega_{M}^{\varepsilon}}I_{M}(w):=\bigcup_{w\in\Omega_{M}^{\varepsilon}}\left(\psi_{M}(w)-\frac{1}{b^{M}},\psi_{M}(w)+\frac{1}{b^{M}}\right).$
We choose some $\varepsilon<\varepsilon(s)$ and we consider integers $K$ large enough s.t. $1/b^{K}<\delta$ and (A.5) holds for all $M\geq K.$ Then,
$\mathcal{H}^{s}_{\delta}(\mathcal{Z}_{b}\cap[0,1))$ $\leq\sum_{M=K}^{\infty}\mathcal{H}^{s}_{\delta}(\{x\,|\,\phi_{M}(x)\in\Omega_{M}^{\varepsilon}\})\leq\sum_{M=K}^{\infty}\sum_{w\in\Omega_{M}^{\varepsilon}}|I_{M}(w)|^{s}$
$\leq\sum_{M=K}^{\infty}|\Omega_{M}^{\varepsilon}|\frac{2^{s}}{b^{sM}}\leq 2^{s}\sum_{M=K}^{\infty}b^{-\frac{s}{2}M}=\frac{2^{s}}{(1-b^{s/2})b^{Ks/2}}.$
The first inequality is a consequence of (A.7) and the fact that the intervals $(I_{M}(w))_{w\in\Omega_{M}^{\varepsilon}}$ cover the set $\{x\,|\,\phi_{M}(x)\in\Omega_{M}^{\varepsilon}\}$. The second line then follows from the choice $\varepsilon<\varepsilon(s),$ and equations (A.5) and (A.2). As we can choose $K$ as large as we want, the last bound reveals $\mathcal{H}^{s}_{\delta}(\mathcal{Z}_{b}\cap[0,1))=0$, which completes the proof.
Acknowledgments
The author would like to thank Verónica Becher and Christoph Aistleitner for their support and for providing valuable references. Special thanks are due to the anonymous referee for numerous helpful comments that substantially improved the paper; in particular, the referee inspired Theorem 1.5 and provided the proof of its second statement.

### v4

@@
-No Nowhere Constant Continuous Function Maps All Non-Normal Numbers to Normal Numbers
+Super-Dense Sets and Their Role in the Theory of Normal Numbers
 Abstract
-In this work, we consider the set of non-normal numbers $\mathcal{N}^{c}$ and ask if there is a non-empty open interval $I$ and a nowhere constant continuous function $\varphi:I\to\mathbb{R}$ which maps all non-normal numbers to normal numbers, i.e., $\varphi(I\cap\mathcal{N}^{c})\subset\mathcal{N}.$ We answer this question negatively.
-We call a set with this property super-dense.
-The super-density of $\mathcal{N}^{c}$ illustrates further its topological size or geometric richness.
-Surprisingly, the measure-theoretically “larger” set of normal numbers $\mathcal{N}$ does not share this property and we give an explicit Cantor-type construction of a nowhere constant continuous function $\hat{C}$, which maps all normal numbers to non-normal numbers.
-Our final result discusses the relation between super-density and the Baire categories employing various set-theoretic strategies.
-Keywords— Normal numbers, Hausdorff dimension, Cantor function
+We introduce and study a new topological notion of the size for subsets of the real line, called super-density.
+A set $A\subset\mathbb{R}$ is super-dense if for every non-empty open interval $I$ and every nowhere constant continuous function $\varphi\colon I\to\mathbb{R}$, we have $\varphi(I\cap A)\cap A\neq\emptyset$.
+We first establish basic properties of super-dense sets.
+Our main topological result characterizes them within the framework of Baire category: a set with the Baire property is super-dense if and only if it is co-meager.
+We then investigate the implications for the theory of normal numbers.
+We prove that the set of non-normal numbers is super-dense, whereas the set of normal numbers is not.
+Consequently, no nowhere constant continuous function can map all non-normal numbers to normal numbers.
+Conversely, we explicitly construct a computable nowhere constant continuous function that maps all normal numbers to non-normal numbers.
+Finally, we provide a constructive algorithm that, given any countable family of nowhere constant continuous functions, produces a real number $x$ such that $x$ and all its images under these functions are non-normal.
+As a corollary, we obtain the existence of a non-normal number $x$ such that $e^{\alpha x}$ is non-normal for every non-zero algebraic $\alpha$.
+Keywords— Normal numbers, Baire categories, Baire sets, Super-density, Hausdorff dimension, Cantor function
 1 Introduction and Main Result
@@
 Although many profound statements about the set of normal numbers have been shown [9, 8, 13, 18, 20, 19, 15] and in recent years there has been increasing interest in the computability of concrete normal numbers [1, 4, 5], the most important conjectures concerning normal numbers remain unsolved.
-In particular, it currently seems unfeasible to show that important mathematical constants like $\sqrt{2},e,$ or $\pi$ are normal.
-In contrast, the field of transcendental number theory has made significant breakthroughs.
+In particular, it currently seems infeasible to show that important mathematical constants like $\sqrt{2},e,$ or $\pi$ are normal.
+By contrast, transcendental number theory has witnessed significant breakthroughs.
 Although many conjectures such as Schanuel’s conjecture regarding transcendental numbers remain open, Hermite, Lindemann and Weierstrass proved in their seminal works the transcendence of $e$ and $\pi$ over a century ago [11, 12, 14, 22].
-This disparity raises a natural question: why is it considerably harder to prove a number is normal than to prove it is transcendental?
-If one understands a proof of normality as a demonstration that a number is not non-normal, the answer seems to be that the set of non-normal numbers is of Hausdorff dimension 1, while the non-transcendental numbers, i.e.
-the algebraic numbers, are only a countable set.
-This note presents another perspective on the complexity of non-normal numbers: There is no nowhere constant continuous function that maps all non-normal numbers to normal numbers.
-This illustrates the intricate geometry of the set of non-normal numbers, potentially explaining the difficulty of proving normality.
-To set the stage, we recall first some standard terminology.
-A real number $x\in{\mathbb{R}}$ can be uniquely expanded with respect to an integer base $b\geq 2$ in the form $x=X_{1}X_{2}\ldots X_{K}.x_{1}x_{2}\cdots$ with digits $X_{i},x_{i}\in\{0,1,\ldots,b-1\}$.
+In fact, the celebrated Lindemann-Weierstrass theorem states that the exponential $e^{\alpha}$ is transcendental for every algebraic number $\alpha\neq 0$.
+Our main results imply that no similar statement holds in the context of (non-)normal numbers.
+Indeed, we construct a non-normal number $x$ such that all exponentials $e^{\alpha x}$ are not normal for every algebraic number $\alpha$.
+Our main insight for the construction of non-normal numbers is to recall that despite being a set of Lebesgue measure zero, the non-normal form a co-meager set and are thus topologically big [2, 16, 17].
+In this work, we introduce the closely related topological concept of super-density which is of independent interest and will allow us to prove non-normal numbers with desired properties.
+First, let us recall the definition of nowhere constant functions:
+Definition 1.1.
+Let $I\subset{\mathbb{R}}$ be a non-empty open interval and $\varphi:I\to{\mathbb{R}}$ some function. We say that $\varphi$ is non-constant at $t_{0}\in I$ if for every $\varepsilon>0$
+$\varphi((t_{0}-\varepsilon,t_{0}+\varepsilon))\setminus\{\varphi(t_{0})\}\neq\emptyset.$ (1.1)
+The function $\varphi$ is nowhere constant (on $I$) if it is non-constant at every $t\in I.$
+We remark that for a function $\varphi$ to be nowhere constant on some interval $I$, it suffices that the set
+$D:=\left\{x\in I\,\bigg|\,\liminf_{h\to 0}\left|\frac{\varphi(x+h)-\varphi(x)}{h}\right|>0\right\}$
+is dense in $I$.
+In particular, a Brownian motion is almost surely (with respect to the canonical Wiener measure) a nowhere constant continuous functions.
+Our main novel concept is content of the following
+Definition 1.2.
+We call a set $A\subset{\mathbb{R}}$ super-dense, if for every non-empty open interval $I$ and every nowhere constant continuous function $\varphi:I\to{\mathbb{R}}$ one has
+$\varphi(I\cap A)\cap A\neq\emptyset.$ (1.2)
+The definition of super-density seems to be not been formulated in the mathematical literature before.
+However, certain constructions are based on related ideas.
+For instance, super-density is vaguely related to the definition of a universally meager set in [21].
+Clearly, a super-dense set is dense in the usual sense with respect to the Euclidean metric; and it is also immediate that a super-dense set is necessarily uncountable.
+Hence, super-dense sets need to be topologically big and the natural question arises to understand the relation between super-density and the well-studied Baire categories.
+We briefly recall that a real subset $A\subset{\mathbb{R}}$ is called meager or of first category if it is given by an at most countable union of nowhere dense sets.
+Otherwise, $A$ is said to be of second category or a fat / non-meager set.
+If the complement $A^{c}:={\mathbb{R}}\setminus A$ is meager, we call $A$ co-meager.
+Finally, a set $A$ has the Baire property if and only if there is an open set $U$ and a meager set $N$ such that $A=U\Delta N$, where $\Delta$ denotes the symmetric difference.
+Our main result below gives a satisfactory relation between super-density and the Baire categories.
+Theorem 1.3.
+The following hold.
+  1. 1.
+    Suppose $A\subset\mathbb{R}$ has the Baire property. Then, $A$ is super-dense if and only if $A$ is co-meager.
+  2. 2.
+    There exists a set $A\subset\mathbb{R}$ such that both $A,A^{c}$ are dense, but not super-dense.
+  3. 3.
+    There exists a set $A\subset\mathbb{R}$ such that both $A,A^{c}$ are super-dense.
+The proof of Theorem 1.3 is spelled out in Section 2.1. A few remarks are in order.
+  1. 1.
+    For general sets $A$ we have the chain of implications $A$ co-meager $\Rightarrow$ AA super-dense $\Rightarrow$ AA non-meager.
+This follows from the observation that co-meager and meager sets always have the Baire property.
+  2. 2.
+    The set $A$ from the third assertion is an example of a super-dense set not being co-meager.
+In view of the first assertion this set $A$ is also an example for a super-dense set without the Baire property.
+  3. 3.
+    By the Baire category theorem, at least one of the sets $A$ or $A^{c}$ from the second assertion is non-meager and, hence, an example for a set of second category which is not super-dense.
+  4. 4.
+    Super-density is in general not stable under intersection, that is, there exist super-dense sets $A,B$ such that $A\cap B$ is not super-dense itself. This is immediate from the third assertion.
+To formulate the implications of Theorem 1.3 for normal numbers, we recall first some standard terminology.
+A real number $x\in{\mathbb{R}}$ has a unique expansion with respect to an integer base $b\geq 2$ in the form $x=X_{1}X_{2}\ldots X_{K}.x_{1}x_{2}\cdots$ with digits $X_{i},x_{i}\in\{0,1,\ldots,b-1\}$.
 This $b$-ary expansion of $x$ is unique if we agree to use a finite expansion, i.e., an expansion with only finitely many nonzero digits, whenever possible.
@@
 Analogously, we introduce the set of all $b$-normal real numbers $\mathcal{N}_{b}$, and the set of all (absolutely) normal numbers $\mathcal{N}$.
-Finally, the complement of a set $A\subset{\mathbb{R}}$ is denoted by $A^{c}:={\mathbb{R}}\setminus A$.
-Our results also need the concept of nowhere constant functions:
-Definition 1.1.
-Let $I\subset{\mathbb{R}}$ be a non-empty open interval and $\varphi:I\to{\mathbb{R}}$ some function. We say that $\varphi$ is non-constant at $t_{0}\in I$ if for every $\varepsilon>0$
-$\varphi((t_{0}-\varepsilon,t_{0}+\varepsilon))\setminus\{\varphi(t_{0})\}\neq\emptyset.$ (1.1)
-The function $\varphi$ is nowhere constant (on $I$) if it is non-constant at every $t\in I.$
-Note that for a function $\varphi$ to be nowhere constant on some interval $I$, it suffices that the set
-$D:=\left\{x\in I\,\bigg|\,\liminf_{h\to 0}\left|\frac{\varphi(x+h)-\varphi(x)}{h}\right|>0\right\}$
-is dense in $I$. In particular, all typical realizations of Brownian motions are nowhere constant continuous functions.
-Our first main result reads as follows:
+Finally, we introduce the set
+$\mathcal{Z}_{b}:=\{x\in{\mathbb{R}}\,|\,\limsup_{M\to\infty}\Lambda_{b,0,M}(x)=1\},$ (1.3)
+which contains all numbers which accumulate zeros in their $b$-adic expansion. Theorem 1.3 implies the following
+Corollary 1.4.
+The following hold.
+  1. 1.
+    The set of normal numbers $\mathcal{N}$ is not super-dense.
+  2. 2.
+    The sets of non-normal numbers $\mathcal{N}^{c}$ and $\mathcal{Z}_{b}$ are super-dense for each integer base $b\geq 2$.
+  3. 3.
+    $\mathcal{Z}_{b}$ has zero Hausdorff dimension for each integer base $b\geq 2$, whereas $\mathcal{N}^{c}$ has full Hausdorff dimension of $1$.
+Proof.
+That $\mathcal{N}^{c}$ has Hausdorff dimension $1$ is well known [10] and the proof that $\mathcal{Z}_{b}$ has zero Hausdorff dimension is presented in the appendix in form of Lemma A.1.
+In view of the first statement Theorem 1.3, it is enough to show that $\mathcal{Z}_{b}$ is co-meager since $\mathcal{N}^{c}$ is then co-meager and super-dense as larger set; and $\mathcal{N}$ is meager, hence not super-dense.
+The following argument is folklore and presented to be self-contained - refined and more subtle analysis can be found in [2, 16, 17].
+We fix an integer base $b\geq 2$ and for $x\in{\mathbb{R}}$ we denote by $x_{i}\in\{0,\ldots,b-1\}$ the decimal places of $x$. We define
+$G_{n}:=\bigcup_{k\geq n}\{x\in{\mathbb{R}}\setminus\mathbb{Q}\,|\,x_{m}=0\text{ for all }m=k,\ldots,k^{2}\}=:\bigcup_{k\geq n}G_{n,k}$
+The set $G_{n}$ is clearly dense. It is also open since each $G_{n,k}$ is a finite union of intervals with rational boundary points. Hence, $G_{n}^{c}$ is nowhere dense and
+$G_{b}:=\bigcap_{n\geq 1}G_{n}$ (1.4)
+is a co-meager $G_{\delta}$ set. Finally we note that $G_{b}\subset\mathcal{Z}_{b}\subset\mathcal{N}_{b,s}^{c}.$
+We see that super-density and the measure-theoretical size of a set are independent of each other.
+Corollary 1.4 says: There is no nowhere constant continuous function that maps all non-normal numbers to normal numbers.
+However, since the proof of Theorem 1.3 is non-constructive, we do not obtain a procedure to compute a non-normal $x$ such that $\varphi(x)$ is also non-normal for a given nowhere constant continuous function $\varphi$.
+The next step - the construction of simultaneously non-normal numbers - is addressed in our next result.
 Theorem 1.5.
-Let $I\subset{\mathbb{R}}$ be any non-empty open interval and $\varphi\colon I\to{\mathbb{R}}$ a nowhere constant continuous function. Then, for every base $b\geq 2$
-$\varphi(I\cap\mathcal{N}_{b,s}^{c})\cap\mathcal{N}_{b,s}^{c}\neq\emptyset.$ (1.2)
-In fact, we have more generally for every countable collection $(\varphi_{k})_{k\in{\mathbb{N}}}$ of nowhere continuous functions $\varphi_{k}\colon I\to{\mathbb{R}}$,
-$\bigcap_{k=1}^{\infty}\varphi_{k}(I\cap\mathcal{N}_{b,s}^{c})\neq\emptyset.$ (1.3)
-Of course, (1.2) directly yields
-$\varphi(I\cap\mathcal{N}^{c})\cap\mathcal{N}^{c}\neq\emptyset,$
-which motivates the title of this note. A few remarks are in order.
+Given a non-empty open interval $I\subset{\mathbb{R}}$, a nowhere constant continuous function $\varphi\colon I\to{\mathbb{R}}$ and integer base $b\geq 2$, there exists an iterative procedure which computes a real number $x\in I$ such that
+$x\in\mathcal{Z}_{b}\,\text{ and }\,\varphi(x)\in\mathcal{Z}_{b}$ (1.5)
+In fact, we have more generally an iterative procedure, which, given a countable collection $(\varphi_{k})_{k\in{\mathbb{N}}}$ of nowhere continuous functions $\varphi_{k}\colon I\to{\mathbb{R}}$, determines a real number $x\in I$ such that
+$\varphi_{k}(x)\in\mathcal{Z}_{b}\text{ for all }k\in\mathcal{Z}_{b}$ (1.6)
+We give a few comments on Theorem 1.5.
   1. 1.
-    The proof of Theorem 1.2 is spelled out in Section 2.1 for a single function $\varphi$ and its full generality in Section 2.3.
-Our method is constructive in the sense that under a few regularity assumptions on the $\varphi_{k}$ we can infer from the proof an algorithm computing inductively the digits of a real number $x$ such that all numbers $\varphi_{k}(x)$ are simultaneously non-normal.
+    The proof of Theorem 1.5 is spelled out in Section 2.2.
+Our method is constructive and the output is computable in the sense that under a few regularity assumptions on the functions $\varphi_{k}$ we can infer from the proof an algorithm computing inductively the digits of a real number $x$ such that all numbers $\varphi_{k}(x)$ are simultaneously non-normal.
 The details are presented in Corollary 2.9 below.
+  2. 2.
     Theorem 1.5 gives a (nearly) optimal result on joint non-normality, which complements the recent work [15], where the construction of simultaneously normal numbers is addressed.
 In the future, we hope to address simultaneous normality and non-normality, that is, to construct for example a normal number $x$ such that $x^{2}$ is not normal - mimicking the conjectured property of $\sqrt{2}$.
-  2. 2.
+  3. 3.
     Let us denote by $\bar{\mathbb{Q}}$ the set of real algebraic numbers.
-Applying Theorem 1.2 to the countable collection $(e^{\alpha x})_{\alpha\in\bar{\mathbb{Q}}\setminus\{0\}}$ and to all open intervals $I=(\alpha,\beta)\subset{\mathbb{R}}$, one concludes that the set of numbers $x$ and all $e^{\alpha x}$ are not normal is a dense subset of $\mathbb{R}$}.
-Note the remarkable difference compared to the case of algebraicity: according to the celebrated Lindemann-Weierstrass Theorem $e^{x}$ is transcendental for every algebraic $x\neq 0$.
-While the existence of such numbers is also a consequence of the set of non-normal numbers being of second category [2, 16, 17], as far as we know there has been no constructive method to generate a non-normal number $x\neq 0$ for which $e^{x}$ is not normal.
-Our proof, however, can easily be turned to an algorithm for the given function $e^{x}$.
-  3. 3.
-    In the setting of a single nowhere constant function $\varphi$, (1.2) can be strengthened further:
-    $\varphi\left(I\cap\bigcap_{b=2}^{\infty}\mathcal{N}_{b,s}^{c}\right)\cap\left(\bigcap_{b=2}^{\infty}\mathcal{N}_{b,s}^{c}\right)\neq\emptyset.$ (1.4)
-    This result can be derived by an adaption of our proof strategy of Theorem 1.2 and can be found below as Corollary 2.10.
+Applying Theorem 1.5 to the countable collection $(e^{\alpha x})_{\alpha\in\bar{\mathbb{Q}}\setminus\{0\}}$ and to all open intervals $I=(\alpha,\beta)\subset{\mathbb{R}}$, one concludes that the set $A:=\{x\,|\,e^{\alpha x}\text{ not normal for all }\alpha\in\bar{\mathbb{Q}}\setminus\{0\}\}$ forms a dense subset of $\mathbb{R}$}.
+This shows that as announced before that there is no analog of the Lindemann-Weierstrass theorem for normal numbers.
+As far as we know, this work is the first to give a constructive method to generate a non-normal number $x\neq 0$ for which $e^{x}$ is not normal.
   4. 4.
-    Equation (1.2) might even hold under the weaker condition that $\varphi$ is just a continuous non-constant function.
-However, our proof fails in this situation and it is unclear to us if the statement of Theorem 1.4 below still holds for non-constant (but not necessarily nowhere constant) functions.
+    Equation 1.6 implies
+    $\bigcap_{k=1}^{\infty}\varphi_{k}(I\cap\mathcal{N}_{b,s}^{c})\neq\emptyset,$ (1.7)
+    which sharpens the super-density of the set $\mathcal{N}_{b,s}^{c}$.
+  5. 5.
+    In the setting of a single nowhere constant function $\varphi$, one can also derive the super-density of the “absolutely non-normal” numbers $\bigcap_{b=2}^{\infty}\mathcal{N}_{b,s}^{c}$.
+The proof is presented in the appendix as Corollary B.1.
+  6. 6.
+    We conjecture that there is in fact no non-constant continuous function $\varphi$ such that $\varphi(\mathcal{N}^{c})\cap\mathcal{N}^{c}=\emptyset$.
+However, our constructive proof fails in this situation and it is unclear to us if the statement of Theorem 1.5 still holds for non-constant (but not necessarily nowhere constant) functions.
 We hope to address this question in future research.
-In our second main result, we want to argue that Theorem 1.2 should not be seen as a reflection of the measure-theoretic size of $\mathcal{N}^{c}$, but rather as a facet of the topological size and the complicated geometry of $\mathcal{N}^{c}$.
-To this end, it is convenient to introduce the following concept.
-Definition 1.3.
-We call a set $A\subset{\mathbb{R}}$ super-dense, if for every non-empty open interval $I$ and every nowhere constant continuous function $\varphi:I\to{\mathbb{R}}$ one has
-$\varphi(I\cap A)\cap A\neq\emptyset.$ (1.5)
-The concept of super-density appears to be novel.
-It is very vaguely related to the definition of a universally meager set, which is a subset $A\subset X$ of a topological space $X$ such that for every Baire space $Y$ and every nowhere constant continuous map $\varphi:Y\to X$ the preimage $\varphi^{-1}(A)$ is meager in $Y$.
-[21].
-Clearly, a super-dense set is dense in the usual sense with respect to the Euclidean metric; and it is not hard to see that a super-dense set is necessarily uncountable.
-More will be said in Theorem 1.5 below.
-Before, we state our second result on (non-)normal numbers.
+The normal numbers $\mathcal{N}$ form the archetype of a set of full measure which is topologically small, i.e., meager and hence not super-dense.
+It is thus of interest to understand if in this case a map $\varphi$ contradicting super-density relies on nonconstructive tools.
+It turns out that we are able to explicitly construct a nowhere constant continuous function with the desired properties.
 Theorem 1.6.
-For every integer $b\geq 2$, there exists a super-dense set $\mathcal{Z}_{b}\subset\mathcal{N}_{b,s}^{c}$ of Hausdorff dimension zero.
-On the other hand, the set of absolutely normal numbers $\mathcal{N}$ is not super-dense.
-The proof of Theorem 1.4 is presented in Section 2.1, where we also give an explicit definition of $\mathcal{Z}_{b}$ in (2.2).
-We note that Theorem 1.4 implies (1.2) (i.e.
-Theorem 1.2 for a single function).
-Due to Borel’s theorem, it might come as a surprise that the non-normal numbers are in some sense “denser” or “richer” than the absolutely normal numbers $\mathcal{N}$, which form a set of full measure.
-Our interpretation is that the geometry of $\mathcal{N}$ is simpler as $\mathcal{N}$ only contains numbers which behave generically, whereas $\mathcal{N}^{c}$ encodes many fractal-like subsets.
-A further intuition for Theorem 1.4 can be found in the fact that $\mathcal{N}$ is a meager set or of first category [2, 16, 17].
-In the following theorem we clarify the relation between super-density and Baire categories.
-Theorem 1.5.
-  1. 1.
-    Let $A\subset\mathbb{R}$ be super-dense. Then, $A$ is also non-meager.
-  2. 2.
-    Let $A\subset\mathbb{R}$ be a co-meager $G_{\delta}$-set, i.e., a countable intersection of open sets. Then, $A$ is super-dense.
-  3. 3.
-    There exists a set $A\subset\mathbb{R}$ such that both $A,A^{c}$ are dense, but not super-dense.
-  4. 4.
-    There exists a set $A\subset\mathbb{R}$ such that both $A,A^{c}$ are super-dense.
-The proof of Theorem 1.5 is spelled out in Section 2.2.
-Note that the first two assertions of Theorem 1.5 directly yield Theorem 1.4 (see also the discussion after Proposition 2.5).
-This more elegant approach has the disadvantage that it is not constructive and does not give an example of a nowhere constant function $\varphi$ such that $\varphi(\mathcal{N})\cap\mathcal{N}=\emptyset$.
-We further observe that due to Baire’s category theorem at least one of $A,A^{c}$ as in the third assertion must be non-meager, dense, but not super-dense.
-Of course, $\mathcal{N}$ is an example for a set which is not super-dense (and meager), and the non-normal numbers $\mathcal{N}^{c}$ are super-dense and non-meager (in fact co-meager).
-Finally, the sets $A,A^{c}$ from 4.
-are super-dense, non-meager, but not co-meager.
+There is an explicit nowhere constant continuous function $\widehat{C}$$\mathbb{R}$$\to{\mathbb{R}}$ such that
+$\widehat{C}(\mathcal{N})\cap\mathcal{N}=\emptyset$ (1.8)
+Moreover, the function $\widehat{C}$$\mathbb{R}$$\to{\mathbb{R}}$ is computable.
+The function $\widehat{C}$ consists of infinitely many self-similar copies of the well-known Cantor function.
+The explicit definition and a proof of the desired properties will be given in Section 2.3, where we also discuss the precise definition of computablility.
 2 Proof of Main Results
 2.1 Proof of Theorem 1.3
-For the proof of Theorem 1.4, some further notation is useful.
-Let $b\geq 2$ and $x=0.x_{1}x_{2}x_{3}\ldots$ be the $b$-ary expansion of some $x\in[0,1)$.
-For $d\in\{0,1,\ldots,b-1\}$ and $M\in{\mathbb{N}}$, we set
-$\Lambda_{b,d,M}(x):=\frac{1}{M}\sum_{j=1}^{M}\chi_{d}(x_{j})$ (2.1)
-with the characteristic function $\chi_{d}$, that is, $\chi_{d}(d)=1$ and $\chi_{d}(k)=0$ for $k\neq d$.
-In other words, $\Lambda_{b,M}(x)$ is the density of the digit $d$ in the first $M$ digits of the $b$-ary expansion of $x$.
-For general $x\in{\mathbb{R}}$, we simply set $\Lambda_{b,d,M}(x):=\Lambda_{b,d,M}(|x|\,\mathrm{mod}\,1).$ We will be interested in the set $\mathcal{Z}_{b}$ of numbers with accumulations of zeros, that is,
-$\mathcal{Z}_{b}:=\{x\in{\mathbb{R}}\,|\,\limsup_{M\to\infty}\Lambda_{b,0,M}(x)=1\}$ (2.2)
-We demonstrate first that the Hausdorff dimension $\dim_{\mathcal{H}}(\mathcal{Z}_{b})=0$ as claimed in Theorem 1.4.
-Lemma 2.1.
-Let $b\geq 2$ be an integer. Then,
-$\dim_{\mathcal{H}}(\mathcal{Z}_{b})=0.$ (2.3)
-The proof of Lemma 2.1 relies on a standard combinatorial estimate and basic properties of Hausdorff measures.
-As it does not directly follow from Egglestone’s result on the Hausdorff-dimension of sets with specific asymptotic digit frequencies [10], we choose to sketch the complete argument in the appendix.
-Proposition 2.2 below demonstrates that $\mathcal{Z}_{b}$ is super-dense and, thus, establishes the first part of Theorem 1.4.
-Proposition 2.2.
-For every base $b\geq 2$, the set $\mathcal{Z}_{b}$ is super-dense, that is, for every non-empty interval $I\subset{\mathbb{R}}$ and every nowhere constant continuous function $\varphi\colon I\to{\mathbb{R}}$,
-$\varphi(I\cap\mathcal{Z}_{b})\cap\mathcal{Z}_{b}\neq\emptyset.$ (2.4)
-The proof of Proposition 2.2 relies on an iterative “zig-zag” construction of an $x\in\mathcal{Z}_{b}$ such that $y=\varphi(x)\in\mathcal{Z}_{b}$, too.
-Through a two-step update we ensure that the $b$-ary expansion of $x$ and $y$ are dominated by zeros infinitely often.
-Proof.
-We fix the basis $b\geq 2$ and assume without loss of generality that $I\subset(0,1)$.
-For an $x\in(0,1)$, we denote by $x=0.d_{1}d_{2}\ldots$ its $b$-ary expansion in digits $d_{i}$.
-Let us introduce the set $T$,
-$T:=\{x\in(0,1)\,|\,\exists L\in{\mathbb{N}}\text{ s.t. }x=0.d_{1}d_{2}\ldots d_{L}\},$ (2.5)
-of numbers with finite $b$-ary expansion and we set
-$\tau(x):=\inf\{L\in{\mathbb{N}}\,|\,x=0.x_{1}\ldots x_{L}\}$ (2.6)
-for $x\in T$ and otherwise $\tau(x)=\infty$. Our construction a second set
-$T^{\prime}:=\bigcup_{x\in T}\bigcup_{K>\tau(x)^{2}+1}\left\{x+\frac{1}{b^{K}}\right\},$ (2.7)
-where the inner union runs over all integers $K>\tau(x)^{2}+1$.
-A number $y\in T^{\prime}$ in the form $x+b^{-K}$ with $x\in T$ and $K>\tau(x)^{2}+1$ and this decomposition is in fact unique as $K$ is determined by the position of the last non-zero digit (which must an isolated ’1’).
-For a $y=x+\frac{1}{b^{K}}\in T^{\prime}$, we similarly set $\tau^{\prime}(y):=K.$ The function $\tau^{\prime}(y)$ is well-defined by our previous observation.
-One easily sees that both, $T$ and $T^{\prime}$, are dense sets in $(0,1)$.
-The idea behind the definition of $T^{\prime}$ is that the zero digits of some $y=x+\frac{1}{b^{K}}\in T^{\prime}$ between the positions $\tau(x)$ and $K$ are preserved under small enough subtractions.
-We now construct 4 sequences
-$(x_{M}^{(1)})_{M\in{\mathbb{N}}},\,(x_{M}^{(2)})_{M\in{\mathbb{N}}},\quad(y_{M}^{(1)})_{M\in{\mathbb{N}}},\,(y_{M}^{(2)})_{M\in{\mathbb{N}}}$
-and prove that $x:=\lim_{M\to\infty}x_{M}^{(2)}\in\mathcal{Z}_{b},$ $y:=\lim_{M\to\infty}y_{M}^{(2)}\in\mathcal{Z}_{b},$ and $y=\varphi(x).$
-Step 1: Choice of starting points and intervals
-Since $T$ is a dense set, we find some $x_{1}^{(1)}\in T\cap I$ and define $L_{x,1}:=\tau(x_{1}^{(1)}).$ Let $y_{1}^{(1)}:=\varphi(x_{1}^{(1)})$ be the corresponding function value.
-Moreover, we introduce the intervals
-$I_{1,x}=(x_{1}^{(1)},x_{1}^{(1)}+b^{-L^{2}_{x,1}-f})\quad I_{1,y}=(-\infty,\infty),$
-where we choose the integer $f\in{\mathbb{N}}$ big enough s.t. the closure of $I_{1,x}$ is still contained in $I$.
-To construct the two other starting values, we recall that $\varphi$ is nowhere constant, and thus we find some $z\in I_{1,x}$ such that $\varphi(z)\neq\varphi(x_{1}^{(1)}).$ Employing the intermediate value theorem and the density of $T^{\prime}$, we in fact find some $x_{1}^{(2)}\in I_{1,x}$ such that $y_{1}^{(2)}:=\varphi(x_{1}^{(2)})\in T^{\prime}$.
-Note that the digits $x_{1}^{(2)}$ at position $L_{x,1}$ to $L_{x,1}^{2}$-th are zero.
-Step 2: Iterative procedure:
-Suppose that for some $M\in{\mathbb{N}}$, we have already defined the values and intervals
-$x_{M}^{(1)}\in T,\,x_{M}^{(2)}\in I_{M,x},\quad y_{M}^{(1)}=\varphi(x_{M}^{(1)}),\,y_{M}^{(2)}=\varphi(x_{M}^{(2)})\in T^{\prime}\cap I_{M,y}.$
-We describe how to construct the $M+1$-th values and intervals. We set $L_{y,M}:=\tau^{\prime}(y_{M}^{(2)})$ and choose $K_{y,M}>L_{y,M}$ big enough s.t.
-$I_{M+1,y}:=\left(y_{M}^{(2)}-\frac{1}{b^{K_{y,M}}},\,y_{M}^{(2)}+\frac{1}{b^{K_{y,M}}}\right)\subset I_{M,y}.$
-The $b$-ary expansion of any $z\in I_{M+1,y}$ agrees with the one of $y_{M}^{(2)}$ up to the index $L_{y,M}-1$ (here we make use of the stability of $T^{\prime}$ under subtraction due to the additional ’1’).
-By continuity of $\varphi$ we find some $x_{M+1}^{(1)}\in T\cap I_{M,x}$ with $y_{M+1}^{(1)}:=\varphi(x_{M+1}^{(1)})\in I_{M+1,y}.$ We set $L_{x,M+1}:=\tau(x_{M+1}^{(1)})$ and
-$I_{M+1,x}:=(x_{M+1}^{(1)},x_{M+1}^{(1)}+b^{-L^{2}_{x,M+1}-1})\cap I_{M,x}\neq\emptyset$
-[…]
-To finish the proof, we observe that both $x_{M}^{(2)}$ and $y_{M}^{(2)}$ are Cauchy sequences and, thus, posses limits $x\in I$ and $y\in{\mathbb{R}}$.
-By continuity of $\varphi$, one has $y=\varphi(x).$ By construction of $x$ one immediately obtains that
-$\Lambda_{b,0,L_{M,x}^{2}}(x)\geq 1-\frac{1}{L_{M,x}},$
-which implies that $x\in\mathcal{Z}_{b}$ if $L_{M,x}\to\infty.$ If $L_{M,x}$ remains bounded the limit $x$ has only finitely many non-zero digits and thus $x\in\mathcal{Z}_{b},$ too.
-One similarly shows that $y\in\mathcal{Z}_{b},$ which completes the proof.
-∎
-It remains to show that the set of absolutely normal numbers $\mathcal{N}$ is not super-dense.
-In fact, we will construct an even larger set $\mathcal{A}\supset\mathcal{N}$, which is not super-dense.
-We define
-[…]
-and recall that the set of $2$-simply normal numbers is given by
-$\mathcal{N}_{s,2}=\left\{x\in{\mathbb{R}}\,|\,\lim_{M\to\infty}\Lambda_{2,0,M}(x)=\frac{1}{2}\right\}.$
-We set
-$\mathcal{A}:=\mathcal{T}_{1}\cap\mathcal{N}_{s,2}.$ (2.8)
-Of course, $\mathcal{N}\subset\mathcal{A}$. Hence, $\mathcal{A}^{c}$ is a null set. We want to show
-Proposition 2.3.
-The set $\mathcal{A}$ is not super-dense.
-Proof.
-The proof heavily relies on properties of Cantor’s function $C:[0,1)\to[0,1),$ also commonly known as devil’s staircase.
-We recall that $C$ is increasing and continuous.
-For our purposes, the following description of $C$ is handy.
-Let $x=0.x_{1}x_{2}x_{3}\ldots$ be the ternary expansion of $x\in[0,1)$.
-If the ternary expansion of $x$ contains no ’1’, one sets $C(x):=x$.
-Otherwise, there exists a minimal index $m\in{\mathbb{N}}$ s.t.
-$x_{m}=1$ and we set $y=C(x)$ to be the number whose binary(!) expansion $y=0.y_{1}y_{2}y_{3}\cdots$ is given by
-$y_{i}:=\begin{cases}0,&\text{ if }i>m,\\ \min\{x_{i},1\},&\text{ if }i\leq m.\end{cases}$
-Note that $C$ maps all numbers in $\mathcal{T}_{1}$ to values with finite binary expansion. That is,
-$C(\mathcal{A}\cap(0,1))\cap\mathcal{A}=\emptyset.$ (2.9)
-Unfortunately, $C$ is locally constant at the deleted thirds in the iterative construction of the Cantor set.
-The main idea is to enrich $C$ by self-similar copies of itself in each interval such that it becomes a nowhere constant function, which is continuous and still satisfies (2.9).
-To this end, it is convenient to work with the symmetrized version $C_{s}$
-$C_{s}(x):=\begin{cases}C(x),&\text{ if }x\leq\frac{1}{2},\\ C(1-x),&\text{ if }x>\frac{1}{2},\end{cases}$ (2.10)
-which still is continuous and preserves the crucial property of the Cantor function that it maps all numbers with at least a ’1’ in the ternary expansion to a number with finite binary expansion.
-We now define the enriched Cantor function $\hat{C}:[0,1]\to[0,1]$.
-To this end, we distinguish the cases, where $x$ has finitely many ’1’s and infinitely many.
-If $x$ has no ’1’, we set $\hat{C}(x)=C_{s}(x).$ If $x$ has $m$ ’1’s in its ternary expansion, we write $x=0.w_{1}w_{2}\ldots w_{m}\sigma,$ where the finite words $w_{i}$ contain exactly an ’1’ as their last digit - we call such words one-stripped in this proof - and $\sigma$ is a word without any ’1’.
-Then, we set
-$\hat{C}(x):=\left(\sum_{r=1}^{m}2^{-r\sum_{k=1}^{r-1}|w_{r}|}C_{s}(w_{r})\right)+2^{-(m+1)\sum_{k=1}^{r}|w_{r}|}C_{s}(\sigma),$ (2.11)
-where the empty sum is understood as zero.
-Here, we abuse notation a bit as we identify a ternary word $w$ with the real number $x=0.w$, i.e.
-$C_{s}(w)$ should be understood as $C_{s}(x)$ with $x=0.w$.
-In the case $m=1$, one applies the normal Cantor function to the first part $w_{1}$, which gives a number terminating after $|w_{1}|$ bits.
-Then, one inserts $|w_{1}|$-many zeros and continues with $\sigma$ - reinterpreted as bit strings.
-For higher $m$, one alternates between inserting an increasing number of zeros and inserting the Cantor bit-string of each $w_{i}$ afterwards.
-If $x$ has infinitely many ’1’s in its ternary expansion, we write $x=0.w_{1}w_{2}\ldots$ with the same definition for $w_{i}$ as above and one similarly sets
-$\hat{C}(x):=\sum_{r=1}^{\infty}2^{-r\sum_{k=1}^{r-1}|w_{r}|}C_{s}(w_{r}).$ (2.12)
-The function $\hat{C}(x)$ is nowhere constant as $C(x)\neq C(y)$ for any $y$ with finite ternary expansion and any $x$ with infinitely many ’0’s and ’2’s.
-Next, we verify that $\hat{C}(x)$ is continuous.
-We consider the approximations $\hat{C}_{k}(x),$ which roughly speaking stop after the $k$-th ’1’ (similarly to the original Cantor function which stops after the first ’1’).
-A rigorous construction goes as follows.
-Let $T_{k}:[0,1)\to[0,1)$ be the truncation map
-[…]
-where the $w_{i}$ are the unique one-stripped words from above with a single ’1’ as their last digit.
-We simply set $\hat{C}_{k}(x):=\hat{C}(T_{k}(x)).$ Let us show that $\hat{C}_{k}(x)$ is continuous.
-We note first that $\hat{C}_{1}(x)=C_{s}(x)$ and for each natural number $k$ the difference $\hat{C}_{k+1}-\hat{C}_{k}$ is nonzero only on closed intervals of the form
-$I_{w_{1},w_{2},\ldots,w_{k}}:=[0.w_{1},w_{2},\ldots,w_{k},w_{1},w_{2},\ldots,w_{k}+3^{-\sum_{j=1}^{k}w_{j}}]$
-with one-stripped $w_{i}$ and on each such interval we have
-$\hat{C}_{k+1}(0.w_{1}\ldots w_{k}\tau)-\hat{C}_{k+1}(0.w_{1}\ldots w_{k}\tau)=2^{-(k+1)\sum_{j=1}^{r}|w_{j}|}C_{s}(\tau)$
-for any ternary word $\tau$.
-The continuity is evident as one inserts a damped copy of the original (symmetrized) Cantor function in each of the separated intervals $I_{w_{1},w_{2},\ldots,w_{k}}$ , where $\hat{C}_{k}$ is constant.
-Moreover, by construction
-$\sup_{x\in(0,1)}|\hat{C}(x)-\hat{C}_{k}(x)|\leq 2^{-k^{2}}$
-so that $\hat{C}_{k}$ converges uniformly to $\hat{C}$. This implies that $\hat{C}$ is continuous, too.
-It remains to show that $\hat{C}(\mathcal{T}_{1})\cap\mathcal{N}_{2,s}=\emptyset.$ To this end, let $x\in\mathcal{T}_{1}$ and we write for its ternary expansion $x=0.w_{1}w_{2}\cdots$ with one-stripped $w_{i}$.
-Let $\hat{w_{i}}$ be the “Cantor image” of $w_{i}$ that is the binary string obtained by flipping each ’2’ in $w_{i}$ to a ’1’.
-Then, the binary expansion of $\hat{C}(x)$ has the form $\hat{C}(x)=\hat{w_{1}}z_{1}\hat{w_{2}}z_{2}\cdots$, where each $z_{r}$ is a string of zeros.
-Now, note that after the last digit of $z_{m}$ the binary expansion of insertion of $\hat{C}(x)$ has at most $\sum_{r=1}^{m}w_{r}$ ’1’s but at least $(m-1)\sum_{r=1}^{m}w_{r}$ zeros.
-That is there a sequence of integers $L_{m}$ s.t.
-$\Lambda_{2,0,L_{m}}(C(x))\geq 1-\frac{1}{m}.$
-This finishes the proof. ∎
-2.2 Proof of Theorem 1.5
-In this section, we prove all statements of Theorem 1.5 one by one. The proofs are all independent of each other and are therefore presented separately. We start with the first statement.
+In this section, we prove all statements of Theorem 1.3 one by one.
+The proofs are all independent of each other and are therefore presented separately.
+The proof of the first assertion requires several preparatory result.
+We start by showing that super dense sets cannot be meager.
 Proposition 2.1.
@@
 $U_{x}:=\{y\in I\,|\,|f(y)-u_{x}|<\varepsilon\}$
-are (relatively) open and let $\hat{U}_{x}$ further be the connected component of $U_{x}$ containing $x$.
-Then, the collection $(\hat{U}_{x})_{x\in I}$ forms an open cover of $I$.
-So, by compactness, there exists a finite cover $\hat{U}_{x_{1}},\ldots\hat{U}_{x_{n}}$ of $I$ with $x_{1}<x_{2}<\cdots<x_{n}$.
-Using again the assumption that $A$ is closed and nowhere dense, we can find open intervals $J_{x_{i},x_{i+1}}\subset(x_{i},x_{i+1}$ such that $J_{x_{i},x_{i+1}}\subset A^{c}$ for $i=1,\ldots,n-1$.
-Finally, we set $I_{x_{i}}:=\hat{U}_{x_{i}}\setminus\left(\bigcup_{j=1}^{n-1}J_{x_{j},x_{j+1}}\right)$ for $i=1,\ldots,n$.
+are (relatively) open and let $\widehat{U}_{x}$ further be the connected component of $U_{x}$ containing $x$.
+Then, the collection $(\widehat{U}_{x})_{x\in I}$ forms an open cover of $I$.
+So, by compactness, there exists a finite cover $\widehat{U}_{x_{1}},\ldots\widehat{U}_{x_{n}}$ of $I$ with $x_{1}<x_{2}<\cdots<x_{n}$.
+Using again the assumption that $A$ is closed and nowhere dense, we can find open intervals $J_{x_{i},x_{i+1}}\subset(x_{i},x_{i+1})$ such that $J_{x_{i},x_{i+1}}\subset A^{c}$ for $i=1,\ldots,n-1$.
+Finally, we set $I_{x_{i}}:=\widehat{U}_{x_{i}}\setminus\left(\bigcup_{j=1}^{n-1}J_{x_{j},x_{j+1}}\right)$ for $i=1,\ldots,n$.
 The intervals $I_{x_{i}},J_{x_{i},x_{i+1}}$ form a partition of $I$ and we define the function $g$ to be $g\equiv u_{x_{i}}$ on $I_{x_{i}}$ and the linear interpolation between $u_{x_{i}}$ and $u_{x_{i+1}}$ on $J_{x_{i},x_{i+1}}$.
@@
 We clearly have $Z\subset\bigcup_{K=1}^{\infty}Z_{K}$ and by the second step each $Z_{K}$ is nowhere dense. Hence, $Z$ is a meager subset of $C(I)$. ∎
+In fact, we can further strengthen the assertion of Proposition 2.1.
+Recall that a set $A\subset{\mathbb{R}}$ is said to be nowhere meager if $A\cap I$ is non-meager for every open interval $I\subset{\mathbb{R}}$.
+Corollary 2.2.
+Let $A\subset\mathbb{R}$ be super-dense. Then, $A$ is nowhere meager.
+Proof.
+We proceed via contradiction.
+Suppose $A\cap I$ is meager for an open interval $I$.
+Super-density is invariant under translation and stretching.
+Hence, we may assume without loss of generality that $B:=A\cap(0,1)$ is meager.
+We set
+$\widehat{B}:=B\cup(1-B)=B\cup\{1-x\,|\,x\in B\}$
+and further define the periodized
+$\mathcal{B}:=\{x+n\,|\,x\in\widehat{B},n\in{\mathbb{Z}}\}.$
+It is easy to see that $\mathcal{B}$ is meager, too.
+By Proposition 2.1, $\mathcal{B}$ is not super-dense and, thus, there exists a nowhere constant continuous function $\varphi:{\mathbb{R}}\to{\mathbb{R}}$ and an open interval $J\subset{\mathbb{R}}$ such that
+$\varphi(J\cap\mathcal{B})\cap\mathcal{B}=\emptyset.$
+Due to the periodicity of $\mathcal{B}$, we may assume that $J\subset(0,1)$. Consider the chainsaw-type function
+$g(x):=\sum_{n\in 2{\mathbb{Z}}}|x-n|\mathbbm{1}_{[-1,1)}(x-n)$
+which is exactly the periodic continuation of the absolute value from [1,1)[-1,1) to $\mathbb{R}$}. We observe that $g$ is continuous and $g\circ\varphi$ is still nowhere constant. Furthermore,
+$(g\circ\varphi)(J\cap\mathcal{B})\cap\mathcal{B}=\emptyset.$
+The reason is that we can write
+$g(x)=\begin{cases}x-2n&\text{ if }x\in[2n,2n+1]\\ 1-(x-2n-1)&\text{ if }x\in[2n+1,2n+2]\end{cases}$
+for $n\in{\mathbb{Z}}$ and the claim follows from the periodicity and reflection invariance of $\mathcal{B}$.
+We further shrink $J$ such that $\{0,1\}\notin(g\circ\varphi)(J)$.
+This is always possible due to continuity and the nowhere constancy of $g\circ\varphi$.
+Since $g\circ\varphi$ takes only values in $(0,1)$ on $J$ we obtain
+$(g\circ\varphi)(J\cap A)\cap A=\emptyset.$
+Here we used that $J\cap A\subset J\cap\mathcal{B}$ for $J\subset(0,1)$. This clearly shows that $A$ cannot be super-dense. ∎
 While Proposition 2.1 gives a rather satisfactory necessary condition for a real set to be super-dense, the next result deals with a complementary sufficient criterion
@@
 Let $A\subset\mathbb{R}$ be a co-meager $G_{\delta}$-set. Then, $A$ is super-dense.
-Let us first demonstrate that the not simply $b$-normal numbers $\mathcal{N}_{b,s}^{c}$ contain a co-meager $G_{\delta}$-set $G_{b}$.
-Then, by Proposition 2.5, $G_{b}$ is super-dense and, thus, $\mathcal{N}_{b,s}^{c}$ is, too.
-As before, for $x\in{\mathbb{R}}$ we denote by $x_{i}\in\{0,\ldots,d-1\}$ the decimal places of $x$.
-We define
-$G_{n}:=\bigcup_{k\geq n}\{x\in{\mathbb{R}}\setminus\mathbb{Q}\,|\,x_{m}=0\text{ for all }m=k,\ldots,k^{2}\}=:\bigcup_{k\geq n}G_{n,k}$
-The set $G_{n}$ is clearly dense. It is also open since each $G_{n,k}$ is a finite union of intervals with rational boundary points. Hence, $G_{n}^{c}$ is nowhere dense and
-$G_{b}:=\bigcap_{n\geq 1}G_{n}$ (2.15)
-is a co-meager $G_{\delta}$ set. Finally we note that $G_{b}\subset\mathcal{Z}_{b}\subset\mathcal{N}_{b,s}^{c}.$
-The following proof follows the suggestion of the anonymous referee.
+The following proof follows the suggestions of the anonymous referee.
 Proof.
 Let $\varphi:\mathbb{R}\to\mathbb{R}$ be continuous and nowhere constant.
-By the intermediate value property, for every open interval $I$ there exists an open subinterval $I_{0}\subseteq I$ such that $\varphi(I_{0})$ contains an open set $J$ in its closure.
+We claim that for every open interval $I$ there exists an open subinterval $I_{0}\subseteq I$ such that $\varphi(I_{0})$ contains an open set $J$ in its closure.
+Indeed, let us fix an open interval $I$ and we pick an $x\in I$.
+Since $I$ is open, we find an $\varepsilon>0$ such $(x-\varepsilon,x+\varepsilon)\subseteq I$.
+Since $\varphi$ is nowhere constant, we find an $y\in(x-\varepsilon,x+\varepsilon)$ such that $\varphi(x)\neq\varphi(y)$.
+We set $I_{0}=(x,y)$.
+We employ the intermediate value theorem to deduce that $\varphi(\overline{I_{0}})\supseteq[\varphi(x),\varphi(y)]$ and the latter clearly contains an open interval since $\varphi(x)\neq\varphi(y)$.
 We now show that $\varphi(A\cap I_{0})$ is co-meager in some open interval.
@@
 $\varphi(A\cap I_{0})\cap A\supseteq(\varphi(A\cap I_{0})\cap J^{\prime})\cap(A\cap J^{\prime})\neq\emptyset.$
+In a next step, we combine our necessary and sufficient criteria from Corollary 2.2 and Proposition 2.3 to deduce the full characterization of super-dense with the Baire property, i.e.
+the first assertion of Theorem 1.3.
+Corollary 2.4.
+Suppose $A\subset{\mathbb{R}}$ has the Baire property. Then, $A$ is super-dense if and only if $A$ is co-meager.
+Proof.
+Suppose $A$ has the Baire property. We show both implications separately from each other.
+Step 1: $A$ co-meager $\Rightarrow$ AA super-dense
+Since $A^{c}$ is meager, there exists a countable family of nowhere dense sets $N_{k}$ such that $A^{c}=\bigcup_{k=1}^{\infty}N_{k}$.
+By definition, the closures $\overline{N}_{k}$ are nowhere dense, too.
+It follows that $\bigcup_{k=1}^{\infty}\overline{N}_{k}$ is a meager set and, hence,
+$G:=\left(\bigcup_{k=1}^{\infty}\overline{N}_{k}\right)^{c}=\bigcap_{k=1}^{\infty}\overline{N}_{k}^{c}$
+is a co-meager $G_{\delta}$-set which is contained in $A$. By Proposition 2.3, $G$ and thus $A$ is super-dense.
+Step 2: $A$ super-dense $\Rightarrow$ AA co-meager
+AA has the Baire property, so there exists an open set $U$ and a meager set $N$ such that $A=U\Delta N.$ We claim that $U$ is a dense open set.
+Indeed, if $U\cap I=\emptyset$ for an open interval, the Baire property implies that $A\cap I$ is meager.
+This contradicts the super-density of $A$ in view of Corollary 2.2.
+But this implies that $A^{c}\subset U^{c}\cup N$, where $U^{c}$ is closed with empty interior, i.e.
+$U^{c}$ is nowhere dense, and by definition $N$ is meager.
+In total, $A^{c}$ is meager.
+∎
 Our proofs of the remaining assertions of Theorem 1.3 are in some sense more explicit.
 We recall the popular Vitali construction of a non-measurable real set $V$: define the equivalence relation on $\mathbb{R}$
-$x\simeq y\Leftarrow x-y\in\mathbb{Q},$
+$x\simeq y\iff x-y\in\mathbb{Q},$
 which naturally decomposes $\mathbb{R}$} into equivalence classes $[x]$.
@@
 It is easy to see that we can assume $V$ to be dense.
-Indeed let $D\subset V$ be some infinitely countable set and $(I_{q_{1},q_{2}})_{q_{1}<q_{2}\in\mathbb{Q}}$ be the countable collection of open intervals with rational endpoints $q_{1},q_{2}$.
-After enumerating both countable sets, we replace each $d_{k}\in D$ by some $d_{k}+r_{k}\in I_{k}$ with $r_{k}\in\mathbb{Q}$.
+Indeed let $D\subset V$ be some infinitely countable set and $(I_{q^{(1)},q^{(2)}})_{q^{(1)}<q^{(2)}\in\mathbb{Q}}$ be the countable collection of open intervals with rational endpoints $q^{(1)},q^{(2)}$.
+Let $d_{1},d_{2},\ldots$ an enumeration of $D$ and similarly we fix an enumeration $I_{q_{k}^{(1)},q_{k}^{(2)}}$ of the rational open intervals.
+In a next step, we replace each $d_{k}\in D$ by some $d_{k}+r_{k}\in I_{q_{k}^{(1)},q_{k}^{(2)}}$ with $r_{k}\in\mathbb{Q}$.
 We fix such dense “Vitali” set $V$ and define $V_{q}:=\{x+q\,|\,x\in V\}$ for $q\in\mathbb{Q}$.
@@
 We denote as usual by $\aleph_{0}$ the cardinality of $\mathbb{N}$} and by the $\aleph_{1}$ the smallest uncountable cardinality, i.e., the cardinality of the set of all countable ordinal numbers.
-The cardinality of the the reals is given by $2^{\aleph_{0}}$ and under the continuum hypothesis $2^{\aleph_{0}}=\aleph_{1}.$ We further recall that any set $X$ can be well-ordered and for such a well-order $\prec$ we set $I_{\prec}(x)=\{t\in X\,|\,t\prec x\}$.
+The cardinality of the reals is given by $2^{\aleph_{0}}$ and under the continuum hypothesis $2^{\aleph_{0}}=\aleph_{1}.$ We further recall that any set $X$ can be well-ordered which we denote by $\prec.$ For a fixed well-order $\prec$ and every $x\in X$, let $I_{\prec}(x)=\{t\in X\,|\,t\prec x\}$.
 Lemma 2.6.
@@
 Having constructed the sets $A^{i}_{f,n}$, we define $A^{i}_{f}=\bigcup_{n\in{\mathbb{N}}}A^{i}_{f,n}$ and by construction these sets have the desired properties.
-We define $A:=\bigcup_{f\in F}A^{1}_{[}f]$ and due to the third property of $A_{f}^{1}$, the set $A$ is super-dense.
+We define $A:=\bigcup_{f\in F}A^{1}_{f}$ and due to the third property of $A_{f}^{1}$, the set $A$ is super-dense.
 Since $A^{c}$ contains all sets $A_{f}^{2}$, it is super-dense by the same reasoning.
@@
 2.2 Proof of Theorem 1.5
+Let us fix some notation. Given an integer base $b\geq 2$, we write $x=0.x_{1}x_{2}x_{3}\ldots$ for the $b$-ary expansion of $x\in[0,1)$. For $d\in\{0,1,\ldots,b-1\}$ and $M\in{\mathbb{N}}$, we set
+$\Lambda_{b,d,M}(x):=\frac{1}{M}\sum_{j=1}^{M}\chi_{d}(x_{j})$ (2.3)
+with the characteristic function $\chi_{d}$, that is, $\chi_{d}(d)=1$ and $\chi_{d}(k)=0$ for $k\neq d$.
+In other words, $\Lambda_{b,d,M}(x)$ is the density of the digit $d$ in the first $M$ digits of the $b$-ary expansion of $x$.
+For general $x\in{\mathbb{R}}$, we simply set $\Lambda_{b,d,M}(x):=\Lambda_{b,d,M}(|x|\,\mathrm{mod}\,1).$ We will be interested in the set $\mathcal{Z}_{b}$ of numbers with accumulations of zeros, that is,
+$\mathcal{Z}_{b}:=\{x\in{\mathbb{R}}\,|\,\limsup_{M\to\infty}\Lambda_{b,0,M}(x)=1\}$ (2.4)
+Proposition 2.8 below gives a constructive demonstration for the super-density of $\mathcal{Z}_{b}$ and, thus, establishes the first part of Theorem 1.5.
+Proposition 2.8.
+Let $b\geq 2$ be an integer base, $I\subset{\mathbb{R}}$ a non-empty open interval, and $\varphi\colon I\to{\mathbb{R}}$ a nowhere constant continuous function.
+Then, there is an iterative algorithm whose outputs converge to a real number $x$ with
+$x\in\mathcal{Z}_{b}\,\text{ and }\,\varphi(x)\in\mathcal{Z}_{b}$ (2.5)
+The proof of Proposition 2.8 relies on an iterative “zig-zag” construction of an $x\in\mathcal{Z}_{b}$ such that $y=\varphi(x)\in\mathcal{Z}_{b}$, too.
+Through a two-step update we ensure that the $b$-ary expansion of $x$ and $y$ are dominated by zeros infinitely often.
+Proof.
+We fix the basis $b\geq 2$ and assume without loss of generality that $I\subset(0,1)$.
+For an $x\in(0,1)$, we denote by $x=0.d_{1}d_{2}\ldots$ its $b$-ary expansion in digits $d_{i}$.
+Let us introduce the set $T$,
+$T:=\{x\in(0,1)\,|\,\exists L\in{\mathbb{N}}\text{ s.t. }x=0.d_{1}d_{2}\ldots d_{L}\},$ (2.6)
+of numbers with finite $b$-ary expansion and we set
+$\tau(x):=\min\{L\in{\mathbb{N}}\,|\,x=0.x_{1}\ldots x_{L}\}$ (2.7)
+for $x\in T$ and otherwise $\tau(x)=\infty$. Our construction requires a second set
+$T^{\prime}:=\bigcup_{x\in T}\bigcup_{K>\tau(x)^{2}+1}\left\{x+\frac{1}{b^{K}}\right\},$ (2.8)
+where the inner union runs over all integers $K>\tau(x)^{2}+1$.
+A number $y\in T^{\prime}$ in the form $x+b^{-K}$ with $x\in T$ and $K>\tau(x)^{2}+1$ and this decomposition is in fact unique as $K$ is determined by the position of the last non-zero digit (which must be an isolated ’1’).
+For a $y=x+\frac{1}{b^{K}}\in T^{\prime}$, we similarly set $\tau^{\prime}(y):=K.$ The function $\tau^{\prime}(y)$ is well-defined by our previous observation.
+One easily sees that both, $T$ and $T^{\prime}$, are dense sets in $(0,1)$.
+The idea behind the definition of $T^{\prime}$ is that the zero digits of some $y=x+\frac{1}{b^{K}}\in T^{\prime}$ between the positions $\tau(x)$ and $K$ are preserved under small enough subtractions.
+We now construct 4 sequences
+$(x_{M}^{(1)})_{M\in{\mathbb{N}}},\,(x_{M}^{(2)})_{M\in{\mathbb{N}}},\quad(y_{M}^{(1)})_{M\in{\mathbb{N}}},\,(y_{M}^{(2)})_{M\in{\mathbb{N}}}$
+and prove that $x:=\lim_{M\to\infty}x_{M}^{(2)}\in\mathcal{Z}_{b},$ $y:=\lim_{M\to\infty}y_{M}^{(2)}\in\mathcal{Z}_{b},$ and $y=\varphi(x).$
+Step 1: Choice of starting points and intervals
+Since $T$ is a dense set, we find some $x_{1}^{(1)}\in T\cap I$ and define $L_{x,1}:=\tau(x_{1}^{(1)}).$ Let $y_{1}^{(1)}:=\varphi(x_{1}^{(1)})$ be the corresponding function value.
+Moreover, we introduce the intervals
+$I_{1,x}=(x_{1}^{(1)},x_{1}^{(1)}+b^{-L^{2}_{x,1}-f})\quad I_{1,y}=(-\infty,\infty),$
+where we choose the integer $f\in{\mathbb{N}}$ big enough s.t. the closure of $I_{1,x}$ is still contained in $I$.
+To construct the two other starting values, we recall that $\varphi$ is nowhere constant, and thus we find some $z\in I_{1,x}$ such that $\varphi(z)\neq\varphi(x_{1}^{(1)}).$ Employing the intermediate value theorem and the density of $T^{\prime}$, we in fact find some $x_{1}^{(2)}\in I_{1,x}$ such that $y_{1}^{(2)}:=\varphi(x_{1}^{(2)})\in T^{\prime}$.
+The digits $x_{1}^{(2)}$ at position $L_{x,1}$ to $L_{x,1}^{2}$-th are zero.
+Step 2: Iterative procedure:
+Suppose that for some $M\in{\mathbb{N}}$, we have already defined the values and intervals
+$x_{M}^{(1)}\in T,\,x_{M}^{(2)}\in I_{M,x},\quad y_{M}^{(1)}=\varphi(x_{M}^{(1)}),\,y_{M}^{(2)}=\varphi(x_{M}^{(2)})\in T^{\prime}\cap I_{M,y}.$
+We describe how to construct the $M+1$-th values and intervals. We set $L_{y,M}:=\tau^{\prime}(y_{M}^{(2)})$ and choose $K_{y,M}>L_{y,M}$ big enough s.t.
+$I_{M+1,y}:=\left(y_{M}^{(2)}-\frac{1}{b^{K_{y,M}}},\,y_{M}^{(2)}+\frac{1}{b^{K_{y,M}}}\right)\subset I_{M,y}.$
+The $b$-ary expansion of any $z\in I_{M+1,y}$ agrees with the one of $y_{M}^{(2)}$ up to the index $L_{y,M}-1$ (here we make use of the stability of $T^{\prime}$ under subtraction due to the additional ’1’).
+By continuity of $\varphi$ we find some $x_{M+1}^{(1)}\in T\cap I_{M,x}$ with $y_{M+1}^{(1)}:=\varphi(x_{M+1}^{(1)})\in I_{M+1,y}.$ We set $L_{x,M+1}:=\tau(x_{M+1}^{(1)})$ and
+$I_{M+1,x}:=(x_{M+1}^{(1)},x_{M+1}^{(1)}+b^{-L^{2}_{x,M+1}-1})\cap I_{M,x}\neq\emptyset.$
+[…]
+To finish the proof, we observe that both $x_{M}^{(2)}$ and $y_{M}^{(2)}$ are Cauchy sequences and, thus, posses limits $x\in I$ and $y\in{\mathbb{R}}$.
+By continuity of $\varphi$, one has $y=\varphi(x).$ By construction of $x$ one immediately obtains that
+$\Lambda_{b,0,L_{M,x}^{2}}(x)\geq 1-\frac{1}{L_{M,x}},$
+which implies that $x\in\mathcal{Z}_{b}$ if $L_{M,x}\to\infty.$ If $L_{M,x}$ remains bounded the limit $x$ has only finitely many non-zero digits and thus $x\in\mathcal{Z}_{b},$ too.
+One similarly shows that $y\in\mathcal{Z}_{b},$ which completes the proof.
+∎
 The proof of Theorem 1.5 in the case of a countable collection of nowhere constant continuous functions $\varphi_{k}:I\to{\mathbb{R}}$ largely builds on the construction for a single function $\varphi$ presented in the proof of Proposition 2.8.
@@
 Here we used the function $\tau^{\prime}$ from the proof of Proposition 2.8.
-Our goal is to find $x^{(m+1)}\in I$ and $y^{(L_{m+1}(\sigma(m+1))}_{\sigma(m+1)}\in T^{\prime}$ such that $|x^{(m+1)}-x^{(m)}|<2^{-(m+2)}$, $\tau^{\prime}\left(y^{(L_{m+1}(\sigma(m+1))}_{\sigma(m+1)}\right)>m$ and (2.16) holds true for $m+1$, too.
+Our goal is to find $x^{(m+1)}\in I$ and $y^{(L_{m+1}(\sigma(m+1)))}_{\sigma(m+1)}\in T^{\prime}$ such that $|x^{(m+1)}-x^{(m)}|<2^{-(m+2)}$, $\tau^{\prime}\left(y^{(L_{m+1}(\sigma(m+1)))}_{\sigma(m+1)}\right)>m$ and (2.9) holds true for $m+1$, too.
 Denote by $U_{m}$ be the open set, for which $z\in U_{m}$ satisfies the condition (2.9) and $|z-x^{(m)}|<2^{-(m+2)}$.
-The standard argument involving the nowhere constancy and continuity of $\varphi_{\sigma(m+1)}$ and the density of $T^{\prime}$, we find an $x^{(m+1)}\in U_{m}$ such that $y^{(L_{m+1}(\sigma(m+1))}_{\sigma(m+1)}:=\varphi_{\sigma(m+1)}(x^{(m+1)})\in T^{\prime}$ with $\tau^{\prime}\left(y^{(L_{m+1}(\sigma(m+1))}_{\sigma(m+1)}\right)>m.$
+The standard argument involving the nowhere constancy and continuity of $\varphi_{\sigma(m+1)}$ and the density of $T^{\prime}$, we find an $x^{(m+1)}\in U_{m}$ such that $y^{(L_{m+1}(\sigma(m+1)))}_{\sigma(m+1)}:=\varphi_{\sigma(m+1)}(x^{(m+1)})\in T^{\prime}$ with $\tau^{\prime}\left(y^{(L_{m+1}(\sigma(m+1)))}_{\sigma(m+1)}\right)>m.$
 Step 3: Conclusion
@@
 Since $x^{(1)}\in(1/4,3/4)$ and $|x^{(m+1)}-x^{(m)}|<2^{-(m+2)}$, we have $x\in(0,1)$.
-Since for any $k\in{\mathbb{N}}$ the sequence $y_{k}^{(m)}$ forms a subsequence of the function values $\varphi_{k}(x_{m})$, we have by continuity $y_{k}:=\varphi_{k}(x)\lim_{m\to\infty}y_{k}^{(m)}$.
+Since for any $k\in{\mathbb{N}}$ the sequence $y_{k}^{(m)}$ forms a subsequence of the function values $\varphi_{k}(x_{m})$, we have by continuity $y_{k}:=\varphi_{k}(x)=\lim_{m\to\infty}y_{k}^{(m)}$.
 On the other hand,
-$|y_{k}^{(m+1)}-y_{k}^{(m)}|<b^{-\tau^{\prime}\left(y_{k}^{(m)}\right)-1)},$
+$|y_{k}^{(m+1)}-y_{k}^{(m)}|<b^{-\tau^{\prime}\left(y_{k}^{(m)}\right)-1},$
 which guarantees that there are infinitely many integers $K\in{\mathbb{N}}$ such that all digits of $y_{k}$ between positions $K$ and $K^{2}$ are zero.
@@
   3. 3.
-    The set $\{\varphi_{k}^{\prime}(x)=0\}$ is finite and its size and elements are computable.
+    The set $\{\varphi_{k}^{\prime}(x)=0\}$ is finite and its number of elements and their values are computable.
 Then, there exists an algorithm computing an $x\in(0,1)$ such that all $\varphi_{k}(x)$ are simultaneously non-normal.
@@
 In contrast to the proof of Theorem 1.5, we cannot set $x^{(m+1)}=w$, but it is enough to find an $z$ such that (2.9) holds for $m+1$.
-We can find such an $x^{(m+1)}$ y a bisection procedure.
+We can find such an $x^{(m+1)}$ by a bisection procedure.
 ∎
-Our last result deals with the non-normality for various bases.
-Corollary 2.10.
-Let $I\subset{\mathbb{R}}$ be any non-empty open interval and $\varphi\colon I\to{\mathbb{R}}$ a nowhere constant continuous function. Then,
-$\varphi\left(I\cap\bigcap_{b=2}^{\infty}\mathcal{N}_{b,s}^{c}\right)\cap\left(\bigcap_{b=2}^{\infty}\mathcal{N}_{b,s}^{c}\right)\neq\emptyset$ (2.17)
+2.3 Proof of Theorem 1.6
+In this section, we construct an explicit function $\widehat{C}$ that contradicts the super-density of $\mathcal{N}$. Let us define
+[…]
+and recall that the set of $2$-simply normal numbers is given by
+$\mathcal{N}_{s,2}=\left\{x\in{\mathbb{R}}\,|\,\lim_{M\to\infty}\Lambda_{2,0,M}(x)=\frac{1}{2}\right\}.$
+We set
+$\mathcal{A}:=\mathcal{T}_{1}\cap\mathcal{N}_{s,2}.$ (2.10)
+Of course, $\mathcal{N}\subset\mathcal{A}$. Hence, $\mathcal{A}^{c}$ is a null set. We want to show
+Proposition 2.10.
+There is an explicit function $\widehat{C}$$\mathbb{R}$$\to{\mathbb{R}}$ such that
+$\widehat{C}(\mathcal{A})\cap\mathcal{A}=\emptyset$ (2.11)
 Proof.
-A constructive proof follows essentially the lines of the proof of Theorem 1.2.
-Instead, we give a topological proof using Theorem 1.5.
-Recall the co-meager $G_{\delta}$-set $G_{b}$ from (2.15).
-We simply define $G:=\bigcup_{b\geq 2}G_{b}$, which is still a co-meager $G_{\delta}$-set.
-By Theorem 1.5, $G$ is super-dense.
-Since $G_{b}\subset\mathcal{Z}_{b}$, we also have $G\subset\bigcap_{b=2}^{\infty}\mathcal{N}_{b,s}^{c}$, which proves the super-density of $\bigcap_{b=2}^{\infty}\mathcal{N}_{b,s}^{c}$.
-∎
-Appendix A Proof of Lemma 2.1
-Proof of Lemma 2.1.
+The proof heavily relies on properties of Cantor’s function $C:[0,1)\to[0,1),$ also commonly known as devil’s staircase.
+We recall that $C$ is increasing and continuous.
+For our purposes, the following description of $C$ is handy.
+Let $x=0.x_{1}x_{2}x_{3}\ldots$ be the ternary expansion of $x\in[0,1)$.
+If the ternary expansion of $x$ contains no ’1’, one sets $C(x):=x$.
+Otherwise, there exists a minimal index $m\in{\mathbb{N}}$ s.t.
+$x_{m}=1$ and we set $y=C(x)$ to be the number whose binary(!) expansion $y=0.y_{1}y_{2}y_{3}\cdots$ is given by
+$y_{i}:=\begin{cases}0,&\text{ if }i>m,\\ \min\{x_{i},1\},&\text{ if }i\leq m.\end{cases}$
+Clearly, $C$ maps all numbers in $\mathcal{T}_{1}$ to values with finite binary expansion. That is,
+$C(\mathcal{A}\cap(0,1))\cap\mathcal{A}=\emptyset.$ (2.12)
+Unfortunately, $C$ is locally constant at the deleted thirds in the iterative construction of the Cantor set.
+The main idea is to enrich $C$ by self-similar copies of itself in each interval such that it becomes a nowhere constant function, which is continuous and still satisfies (2.12).
+To this end, it is convenient to work with the symmetrized version $C_{s}$
+$C_{s}(x):=\begin{cases}C(x),&\text{ if }x\leq\frac{1}{2},\\ C(1-x),&\text{ if }x>\frac{1}{2},\end{cases}$ (2.13)
+which still is continuous and preserves the crucial property of the Cantor function that it maps all numbers with at least a ’1’ in the ternary expansion to a number with finite binary expansion.
+We now define the enriched Cantor function $\widehat{C}:[0,1]\to[0,1]$.
+To this end, we distinguish the cases, where $x$ has finitely many ’1’s and infinitely many.
+If $x$ has no ’1’, we set $\widehat{C}(x)=C_{s}(x).$ If $x$ has $m$ ’1’s in its ternary expansion, we write $x=0.w_{1}w_{2}\ldots w_{m}\sigma,$ where the finite words $w_{i}$ contain exactly an ’1’ as their last digit - we call such words one-stripped in this proof - and $\sigma$ is a word without any ’1’.
+Then, we set
+$\widehat{C}(x):=\left(\sum_{r=1}^{m}2^{-r\sum_{k=1}^{r-1}|w_{r}|}C_{s}(w_{r})\right)+2^{-(m+1)\sum_{k=1}^{r}|w_{r}|}C_{s}(\sigma),$ (2.14)
+where the empty sum is understood as zero.
+Here, we abuse notation a bit as we identify a ternary word $w$ with the real number $x=0.w$, i.e.
+$C_{s}(w)$ should be understood as $C_{s}(x)$ with $x=0.w$.
+In the case $m=1$, one applies the normal Cantor function to the first part $w_{1}$, which gives a number terminating after $|w_{1}|$ bits.
+Then, one inserts $|w_{1}|$-many zeros and continues with $\sigma$ - reinterpreted as bit strings.
+For higher $m$, one alternates between inserting an increasing number of zeros and inserting the Cantor bit-string of each $w_{i}$ afterwards.
+If $x$ has infinitely many ’1’s in its ternary expansion, we write $x=0.w_{1}w_{2}\ldots$ with the same definition for $w_{i}$ as above and one similarly sets
+$\widehat{C}(x):=\sum_{r=1}^{\infty}2^{-r\sum_{k=1}^{r-1}|w_{r}|}C_{s}(w_{r}).$ (2.15)
+The function $\widehat{C}(x)$ is nowhere constant as $C(x)\neq C(y)$ for any $y$ with finite ternary expansion and any $x$ with infinitely many ’0’s and ’2’s.
+Next, we verify that $\widehat{C}(x)$ is continuous.
+We consider the approximations $\widehat{C}_{k}(x),$ which roughly speaking stop after the $k$-th ’1’ (similarly to the original Cantor function which stops after the first ’1’).
+A rigorous construction goes as follows.
+Let $T_{k}:[0,1)\to[0,1)$ be the truncation map
+[…]
+where the $w_{i}$ are the unique one-stripped words from above with a single ’1’ as their last digit.
+We simply set $\widehat{C}_{k}(x):=\widehat{C}(T_{k}(x)).$ Let us show that $\widehat{C}_{k}(x)$ is continuous.
+We note first that $\widehat{C}_{1}(x)=C_{s}(x)$ and for each natural number $k$ the difference $\widehat{C}_{k+1}-\widehat{C}_{k}$ is nonzero only on closed intervals of the form
+$I_{w_{1},w_{2},\ldots,w_{k}}:=[0.w_{1},w_{2},\ldots,w_{k},w_{1},w_{2},\ldots,w_{k}+3^{-\sum_{j=1}^{k}w_{j}}]$
+with one-stripped $w_{i}$ and on each such interval we have
+$\widehat{C}_{k+1}(0.w_{1}\ldots w_{k}\tau)-\widehat{C}_{k}(0.w_{1}\ldots w_{k}\tau)=2^{-(k+1)\sum_{j=1}^{k}|w_{j}|}C_{s}(\tau)$
+for any ternary word $\tau$.
+The continuity is evident as one inserts a damped copy of the original (symmetrized) Cantor function in each of the separated intervals $I_{w_{1},w_{2},\ldots,w_{k}}$ , where $\widehat{C}_{k}$ is constant.
+Moreover, by construction
+$\sup_{x\in(0,1)}|\widehat{C}(x)-\widehat{C}_{k}(x)|\leq 2^{-k^{2}}$ (2.16)
+so that $\widehat{C}_{k}$ converges uniformly to $\widehat{C}$. This implies that $\widehat{C}$ is continuous, too.
+It remains to show that $\widehat{C}(\mathcal{T}_{1})\cap\mathcal{N}_{2,s}=\emptyset.$ To this end, let $x\in\mathcal{T}_{1}$ and we write for its ternary expansion $x=0.w_{1}w_{2}\cdots$ with one-stripped $w_{i}$.
+Let $\widehat{w_{i}}$ be the “Cantor image” of $w_{i}$ that is the binary string obtained by flipping each ’2’ in $w_{i}$ to a ’1’.
+Then, the binary expansion of $\widehat{C}(x)$ has the form $\widehat{C}(x)=\widehat{w_{1}}z_{1}\widehat{w_{2}}z_{2}\cdots$, where each $z_{r}$ is a string of zeros.
+Now, observe that after the last digit of $z_{m}$ the binary expansion of insertion of $\widehat{C}(x)$ has at most $\sum_{r=1}^{m}w_{r}$ ’1’s but at least $(m-1)\sum_{r=1}^{m}w_{r}$ zeros.
+That is there a sequence of integers $L_{m}$ s.t.
+$\Lambda_{2,0,L_{m}}(\widehat{C}(x))\geq 1-\frac{1}{m}.$
+This shows that $\widehat{C}$ has all desired properties on $[0,1]$. Since $\widehat{C}(0)=\widehat{C}(1)=0$, the periodic extension of $\widehat{C}$ to the whole real line is the desired function. ∎
+In a next step, we want to show that the function $\widehat{C}$ is in fact computable.
+We follow the conventions from [23] where the reader may find more details.
+A name of a real number $x\in{\mathbb{R}}$ is a sequence of rational numbers $p_{n}$ with $|x-p_{n}|<2^{-n}$.
+A real function $f:I\to{\mathbb{R}}$ is computable if and only if there is a (Type-2) machine which transforms any name of $x\in I$ to a name of $f(x)$.
+Computability is a restrictive property as it requires the function to be continuous.
+We recall some key properties.
+  1. 1.
+    If $f,g$ are computable functions and $a$ is a computable real functions, then $af(x)+g(x),f(x)\cdot g(x),1/f(x)$ and $\min\{f(x),g(x)\}$ are also computable functions [23, Corollary 4.3.4].
+  2. 2.
+    The composition $f\circ g$ of two computable functions $f,g$ is computable (if well-defined) [23, Lemma 2.3.18].
+  3. 3.
+    Let $F:=\sum_{k=j}^{\infty}f_{j}$.
+Suppose that each $f_{j}$ is computable and that there exists a computable sequence of natural numbers $M_{k}$ such that $\sum_{j=M_{k}}^{\infty}\|f_{j}\|_{\infty}<2^{-k}$.
+Then, $F$ is computable [23, Theorem 4.3.8].
+  4. 4.
+    Suppose $f:[0,1]\to{\mathbb{R}}$ is computable and $f(0)=f(1)=0.$ The periodic continuation of $f$ is also computable. This follows for instance from [23, Lemma 4.3.5]
+Theorem 1.6 follows from the following
+Proposition 2.11.
+The function $\widehat{C}$ from Proposition 2.10 is computable.
+Proof.
+Step 1: The Cantor function $C:[0,1]\to[0,1]$ is computable.
+This fact is well-known, but hard to trace in the mathematical literature. We give a sketch of the proof. We construct a uniform Type–2 machine realizing $C$.
+Let $\rho$ denote the standard Cauchy representation of $\mathbb{R}$ [23, Definition 4.1.5].
+Thus a $\rho$–name of $x\in[0,1]$ is a sequence $(p_{n})_{n\in\mathbb{N}}$ of rationals satisfying $|p_{n}-x|\leq 2^{-n}$ for all $n$.
+Define a partial operator
+$M\colon\subseteq\mathbb{N}^{\mathbb{N}}\to\mathbb{N}^{\mathbb{N}}$
+as follows. On input a $\rho$–name $p=(p_{n})_{n}$ of $x$ and on input $k\in\mathbb{N}$, the machine $M$ proceeds uniformly:
+  1. 1.
+    Choose the least $n=n(k)$ such that $2^{-n}<3^{-(k+1)}$.
+  2. 2.
+    Using $p_{n}$, determine the unique integer $m\in\{0,\dots,3^{k}-1\}$ with
+    $x\in\left[\frac{m}{3^{k}},\frac{m+1}{3^{k}}\right].$
+    This uniquely fixes the first $k$ digits of the ternary expansion of $x$.
+  3. 3.
+    Replace each ternary digit 0 by 0 and each ternary digit $2$ by $1$. If a digit $1$ occurs, the resulting binary expansion is finite.
+  4. 4.
+    Let $q_{k}$ be the dyadic rational determined by the resulting $k$ binary digits and output $M(p)(k):=q_{k}$.
+By construction,
+$|q_{k}-C(x)|\leq 2^{-k}$
+for all $k\in\mathbb{N}$. Hence $(q_{k})_{k}$ is a $\rho$–name of $C(x)$. The procedure is uniform and independent of the particular input $x$. Therefore $M$ is a realizer of $C$, i.e.
+$\rho(M(p))=C(\rho(p))$
+for all $\rho$–names $p$ of $x\in[0,1]$, and $C$ is $\rho$–computable.
+Step 2: The function $\widehat{C}$$\mathbb{R}$$\to{\mathbb{R}}$ is computable.
+This turns out to be largely a consequence of the above discussed properties of computability.
+Since $\widehat{C}$ is defined on $\mathbb{R}$} via a periodic extension, it is enough to show that $\widehat{C}:[0,1]\to{\mathbb{R}}$ is computable.
+Moreover, employing the functions $\widehat{C}_{k}$ from the proof of Proposition 2.10, we can write
+$\widehat{C}(x)=\sum_{k=1}^{\infty}\widehat{C}_{k}(x)-\widehat{C}_{k-1}(x),$
+where $\widehat{C}_{0}\equiv 0.$ In view of (2.16), it is enough to show that each difference $\widehat{C}_{k}-\widehat{C}_{k-1}$ is computable.
+For $k=1$, this follows from the representation $C_{s}(x)=\min\{1-C(x),C(x)\}$ and the computability of the Cantor function.
+For $k>1$, we use the series representation
+$\widehat{C}_{k}(x)-\widehat{C}_{k-1}(x):=\sum_{w_{1},\ldots,w_{k-1}\text{ one-stripped}}\mathbbm{1}_{x=0.w_{1}\ldots w_{k-1}\tau}2^{-k\sum_{j=1}^{k-1}|w_{j}|}C_{s}(\tau).$
+Note that each summand is computable as rescaled Cantor function. Due to the explicit damping, the full series is computable, too. This completes the proof. ∎
+Appendix A Proof of $\dim_{\mathcal{H}}(\mathcal{Z}_{b})=0$
+The main purpose of this appendix is to present a proof of the following
+Lemma A.1.
+Let $b\geq 2$ be an integer. Then,
+$\dim_{\mathcal{H}}(\mathcal{Z}_{b})=0.$ (A.1)
+The proof of Lemma A.1 relies on a standard combinatorial estimate and basic properties of Hausdorff measures.
+As it does not directly follow from Egglestone’s result on the Hausdorff-dimension of sets with specific asymptotic digit frequencies [10], we choose to sketch the complete argument.
+Proof of Lemma A.1.
 First, we recall the construction of the (outer) Hausdorff measure on the real line. For real numbers $s,\delta>0$ we introduce the functional
@@
 $\Omega_{M}^{\varepsilon}:=\{w\in\Omega_{M}\,|\,N_{m}(w)\geq 1-\varepsilon\}.$
-We note that the size of $\Omega_{M}^{\varepsilon}$ can be written in terms of a binomial expansion
-$|\Omega_{M}^{\varepsilon}|=\sum_{K\geq(1-\varepsilon)M}^{M}\binom{M}{K}(b-1)^{M-K}\leq\varepsilon M\binom{M}{\lceil(1-\varepsilon)M\rceil}(1-b)^{\varepsilon N},$ (A.3)
+The size of $\Omega_{M}^{\varepsilon}$ can be written in terms of a binomial expansion
+$|\Omega_{M}^{\varepsilon}|=\sum_{K\geq(1-\varepsilon)M}^{M}\binom{M}{K}(b-1)^{M-K}\leq\varepsilon M\binom{M}{\lceil(1-\varepsilon)M\rceil}(1-b)^{\varepsilon M},$ (A.4)
 where the bound follows from the monotonicity of the binomial. We now employ Stirling’s formula to obtain the asymptotic expansion
@@
 $\mathcal{Z}_{b}\cap[0,1)\subset\bigcup_{M=K}^{\infty}\{x\,|\,N_{M}(x)\geq 1-\varepsilon\}=\bigcup_{M=K}^{\infty}\{x\,|\,\phi_{M}(x)\in\Omega_{M}^{\varepsilon}\}$ (A.8)
-for any integer $K$ and any $\varepsilon>0.$ Moreover, we note that
+for any integer $K$ and any $\varepsilon>0.$ Moreover,
 $\{x\,|\,\phi_{M}(x)\in\Omega_{M}^{\varepsilon}\}\subset\bigcup_{w\in\Omega_{M}^{\varepsilon}}I_{M}(w):=\bigcup_{w\in\Omega_{M}^{\varepsilon}}\left(\psi_{M}(w)-\frac{1}{b^{M}},\psi_{M}(w)+\frac{1}{b^{M}}\right).$
@@
 As we can choose $K$ as large as we want, the last bound reveals $\mathcal{H}^{s}_{\delta}(\mathcal{Z}_{b}\cap[0,1))=0$, which completes the proof.
+Appendix B Simultaneous non-normality for different bases
+In this appendix, we discuss the non-normality for various bases.
+Corollary B.1.
+Let $I\subset{\mathbb{R}}$ be any non-empty open interval and $\varphi\colon I\to{\mathbb{R}}$ a nowhere constant continuous function. Then,
+$\varphi\left(I\cap\bigcap_{b=2}^{\infty}\mathcal{N}_{b,s}^{c}\right)\cap\left(\bigcap_{b=2}^{\infty}\mathcal{N}_{b,s}^{c}\right)\neq\emptyset$ (B.1)
+Proof.
+A constructive proof follows essentially the lines of the proof of Theorem 1.5.
+Instead, we give a topological proof using Theorem 1.3.
+Recall the co-meager $G_{\delta}$-set $G_{b}$ from (1.4).
+We simply define $G:=\bigcup_{b\geq 2}G_{b}$, which is still a co-meager $G_{\delta}$-set.
+By Theorem 1.3, $G$ is super-dense.
+Since $G_{b}\subset\mathcal{Z}_{b}$, we also have $G\subset\bigcap_{b=2}^{\infty}\mathcal{N}_{b,s}^{c}$, which proves the super-density of $\bigcap_{b=2}^{\infty}\mathcal{N}_{b,s}^{c}$.
+∎
 Acknowledgments
 The author would like to thank Verónica Becher and Christoph Aistleitner for their support and for providing valuable references.
-Special thanks are due to the anonymous referee for numerous helpful comments that substantially improved the paper; in particular, the referee inspired Theorem 1.5 and provided the proof of its second statement.
+Special thanks are due to the anonymous referee for numerous helpful comments on the structure and the content of this work that substantially improved the paper.
+In particular, the referee inspired Theorem 1.3 and provided proof ideas for its first statement.
+More precisely, the proofs of Proposition 2.3 and Corollary 2.4 follow his suggestions.